### v2

# Event-Based Control via Sparsity-Promoting Regularization: A Rollout Approach with Performance Guarantees

Abstract
This paper presents a controller design framework aiming to balance control performance and actuation rate. Control performance is evaluated by an infinite-horizon average cost, and the number of control actions is penalized via sparsity-promoting regularization. Since the formulated optimal control problem has a combinatorial nature, we employ a rollout algorithm to obtain a tractable suboptimal solution. In the proposed scheme, actuation timings are determined through a multistage minimization procedure based on a receding-horizon approach, and the corresponding control inputs are computed online. We establish theoretical performance guarantees with respect to periodic control and prove the stability of the closed-loop system. The effectiveness of the proposed method is demonstrated through a numerical example.
I Introduction
Sparse control, in which the control signal remains zero for as long as possible within a control horizon, has attracted much attention in systems and control. Such control methods facilitate achieving control objectives with limited resource usage in networked control systems, since sparse signals admit compact representations [1]. Moreover, sparse control has the potential to reduce power consumption by saving actuation durations, which is crucial in applications such as railways [2] and electric vehicles [3]. These advantages have motivated development on the theoretical foundations of sparse control, including controllability [4], stability [5], and optimal control [6, 7]. A central challenge lies in developing frameworks that achieve desirable control performance with sparse inputs, due to the fundamental trade-off between closed-loop performance and input sparsity [8].
Sparsity of the control signal can be characterized in various ways, such as by the duration of nonzero inputs in an element-wise sense or by the number of active actuators at each time step. Among these notions of sparsity, this paper focuses on reducing the actuation rate, which gives rise to intermittent actuation, where the control input remains identically zero over consecutive time intervals. Such sparsity arises in practice; for example, in energy packet networks [9], where an on-demand power supply delivers energy to actuators intermittently. Motivated by this practical relevance, we investigate a sparse control method that achieves desirable control performance under intermittent actuation.
In line with this direction, controller design methods for achieving desired performance under intermittent actuation have been investigated. For example, [10, 11, 12] have examined optimal control problems under constraints on the number of control actions. Alternatively, [13] and [14] have incorporated sparsity-promoting regularization terms into the optimization problem instead of enforcing explicit constraints. These studies provide near-optimal actuation strategies with theoretical performance bounds; however, their problem settings are restricted, such as considering noise-free systems [14, 12] or excluding control energy from the objective function [10, 11, 13]. For the case where the objective function consists of the state trajectory and the control input, [15] studied quadratic cost minimization for noise-affected systems under a threshold-based event-triggering rule combined with packetized dead-beat control. While their work quantifies the trade-off between communication and control performance, the actuation timings are determined by the predesigned threshold rule rather than being optimized. This challenge highlights the need for a principled method that jointly optimizes discrete actuation timings and continuous control laws to balance control performance and actuation rate, while providing provable performance improvements.
One promising approach to such co-design problems is the rollout algorithm, a sequential optimization technique in dynamic programming [16, 17]. In a rollout scheme, control laws and scheduling decisions over a finite horizon are optimized sequentially. For example, [18] has addressed quadratic cost minimization under an average communication rate constraint and developed a rollout-based event-triggered control method that provides both performance and stability guarantees. This method achieves strictly better performance than the periodic control strategy under the same communication rate; however, its flexibility is limited by the assumption of a uniform communication rate within each horizon. Furthermore, [19] has studied the simultaneous minimization of a finite-horizon quadratic cost and the number of control actions. The authors proposed an algorithm that sequentially minimizes the one-step cost involving the control input and actuation timing, providing theoretical performance guarantees but no stability results.
In this paper, we propose a controller design method that aims to balance control performance and actuation rate, which provides both performance and stability guarantees. To promote sparsity in the control signal, we introduce an actuation-rate regularization term into the optimization problem, instead of imposing a constraint on the number of control actions. We formulate the problem as a co-design of control inputs and actuation timings, and seek to find control and triggering policies that minimize an objective function combining the LQ cost with a sparsity-promoting term. Based on the rollout strategy, control and triggering policies over multiple stages are optimized, assuming that periodic control, where the input is applied at fixed time instants, is used beyond the horizon. This procedure provides a more general framework than that in [19]. We show that the proposed method achieves theoretical performance bounds with respect to periodic control and ensures mean-square stability of the closed-loop system. While our techniques in performance analysis are similar to those in [18], it is not straightforward to apply their results to the present work, since the trade-off between control performance and communication rate is formulated in a different way. Moreover, whereas [18] establishes stochastic stability of the system in terms of ergodicity, obtained as a byproduct of the performance analysis, this stability notion cannot directly extend to the output-feedback setting considered in this paper. A numerical example demonstrates the effectiveness of the proposed algorithm.
The remainder of this paper is organized as follows. In section II, we describe the problem formulation. Section III provides some preliminaries and presents the proposed algorithm, followed by a theoretical analysis of our framework. Section V illustrates a numerical example, and Section VI concludes this paper.
Notation
We denote $\mathbb{R}$ the set of real numbers, $\mathbb{N}$ the set of natural numbers, and $\mathbb{Z}_{\geq 0}$ the set of nonnegative integers. The $n\times n$ identity matrix is denoted by $I_{n}$. For a positive-semidefinite matrix $P$, its square root matrix is expressed by $P^{1/2}$. For a random variable $x$, $\mathbb{P}(x)$ is the probability of $x$, and $\mathbb{E}[x]$ and $\mathbb{E}\left[x\,\middle|\,\cdot\right]$ represent the expectation and the conditional expectation, respectively. For integers $p$ and $q$, the notation $p\equiv 0\pmod{q}$ indicates that $p$ is a multiple of $q$. For a set $\mathcal{A}$, the indicator function is defined as
$\mathbb{I}_{\mathcal{A}}(x)\coloneqq\begin{cases}1&\text{if }x\in\mathcal{A},\\ 0&\text{if }x\notin\mathcal{A}.\end{cases}$ (1)
II Problem Formulation
We consider the following discrete-time linear system:
$x_{k+1}$ $=Ax_{k}+Bu_{k}+w_{k},$ (2)
$y_{k}$ $=Cx_{k}+v_{k},$ (3)
where $x_{k}\in\mathbb{R}^{n_{x}}$, $u_{k}\in\mathbb{R}^{n_{u}}$, and $y_{k}\in\mathbb{R}^{n_{y}}$ are the state, the control input, and the output, respectively. The random signals $w_{k}$ and $v_{k}$ are assumed to be independent and identically distributed (i.i.d.) Gaussian processes with mean zero and covariance matrices $\Omega_{w}\succ 0$ and $\Omega_{v}\succ 0$, respectively. The initial state $x_{0}$ follows a Gaussian distribution with mean $\bar{x}_{0}$ and covariance matrix $\Omega_{x_{0}}\succeq 0$, and is statistically uncorrelated with both $w_{k}$ and $v_{k}$ for all $k\in\mathbb{Z}_{\geq 0}$. The pairs $(A,B)$ and $(A,\Omega_{w}^{1/2})$ are assumed to be controllable, and $(A,C)$ is assumed to be observable.
[…]
Figure 1: Considered feedback system.
The block diagram of the considered system is depicted in Fig. 1. The controller transmits control signals to the actuators only at selected time instants, while the control input is set to zero at all other time steps. We represent this actuation strategy by a binary variable $\delta_{k}\in\{0,1\}$, such that $u_{k}=0$ if $\delta_{k}=0$, and $u_{k}\in\mathbb{R}^{n_{u}}$ otherwise. At each time $k$, the controller determines $u_{k}$ and $\delta_{k}$ based on the information available up to time $k$. The information set is formally defined as
$\mathcal{I}_{k}$ $\coloneqq\{y_{i},u_{j},\delta_{j}\colon 0\leq i\leq k,\ 0\leq j\leq k-1\},$ (4)
$\mathcal{I}_{0}$ $\coloneqq\{y_{0}\}.$ (5)
Let $\mu^{u}_{k}\colon\mathcal{I}_{k}\to\mathbb{R}^{n_{u}}$ and $\mu^{\delta}_{k}\colon\mathcal{I}_{k}\to\{0,1\}$ denote the control policy and the triggering policy, respectively, both of which are $\sigma(\mathcal{I}_{k})$-measurable. Then, the control input $u_{k}$ and the triggering variable $\delta_{k}$ are described as
$u_{k}=\mu^{u}_{k}(\mathcal{I}_{k}),\quad\delta_{k}=\mu^{\delta}_{k}(\mathcal{I}_{k}).$ (6)
We denote the sequences of these policies by $\mu^{u}\coloneqq\{\mu^{u}_{k}\}_{k\in\mathbb{Z}_{\geq 0}}$ and $\mu^{\delta}\coloneqq\{\mu^{\delta}_{k}\}_{k\in\mathbb{Z}_{\geq 0}}$, respectively.
At each time step, the controller computes a state estimation based on the available information. The estimation $\hat{x}_{k}\coloneqq\mathbb{E}[x_{k}\mid\mathcal{I}_{k}]$ [^1] Throughout this paper, $\mathbb{E}[\cdot\mid\mathcal{I}_{k}]$ denotes shorthand for $\mathbb{E}[\cdot\mid\sigma(\mathcal{I}_{k})]$. is updated by
$\hat{x}_{k+1}=A\hat{x}_{k}+Bu_{k}+G_{k+1}(y_{k+1}-C(A\hat{x}_{k}+Bu_{k}))$ (7)
with the initial condition $\hat{x}_{0}=\bar{x}_{0}+G_{0}(y_{0}-C\bar{x}_{0})$, where $G_{k}=\Sigma^{-}_{k}C^{\top}(C\Sigma^{-}_{k}C^{\top}+\Omega_{v})^{-1}$ is the Kalman gain. The matrix $\Sigma^{-}_{k}$ is given by
$\Sigma^{-}_{k}$ $=A\Sigma_{k}A^{\top}+\Omega_{w},$ (8)
$\Sigma_{k}$ $=\Sigma^{-}_{k}-\Sigma^{-}_{k}C^{\top}\left(C\Sigma^{-}_{k}C^{\top}+\Omega_{v}\right)^{-1}C\Sigma^{-}_{k},$ (9)
where $\Sigma^{-}_{0}=\Omega_{x_{0}}$ and $\Sigma_{k}=\mathbb{E}[(x_{k}-\hat{x}_{k})(x_{k}-\hat{x}_{k})^{\top}\mid\mathcal{I}_{k}]$. By the controllability and the observability assumptions, there exist positive-definite matrices $\tilde{\Sigma}$ and $\Sigma$ such that
$\lim_{k\to\infty}\Sigma^{-}_{k}=\tilde{\Sigma},\quad\lim_{k\to\infty}\Sigma_{k}=\Sigma,$ (10)
and the steady-state Kalman gain is given by
$G=\tilde{\Sigma}C^{\top}\left(C\tilde{\Sigma}C^{\top}+\Omega_{v}\right)^{-1}.$ (11)
The objective of this paper is to design $\mu^{u}$ and $\mu^{\delta}$ so as to balance the trade-off between control performance and the number of control actions. The average control performance is measured by the following quadratic cost:
$J^{a}_{c}(\mu^{u},\mu^{\delta})=\limsup_{N\to\infty}\frac{1}{N}\mathbb{E}\left[\sum_{k=0}^{N-1}x_{k}^{\top}Qx_{k}+u_{k}^{\top}Ru_{k}\right],$ (12)
where $Q\succeq 0$ and $R\succ 0$. We assume that the pair $(A,Q^{1/2})$ is observable. To capture the sparsity of the control input, we introduce the average actuation rate as
$J^{a}_{r}(\mu^{\delta})=\limsup_{N\to\infty}\frac{1}{N}\mathbb{E}\left[\sum_{k=0}^{N-1}\delta_{k}\right].$ (13)
Using (12) and (13), we formulate the co-design problem as
$\underset{\displaystyle\mu^{u},\mu^{\delta}}{\mathrm{min}}\quad J^{a}(\mu^{u},\mu^{\delta})=J^{a}_{c}(\mu^{u},\mu^{\delta})+\theta J^{a}_{r}(\mu^{\delta})\hfil\hfil\hfil\hfil$ (14)
[…]
where $\theta>0$ is a weighting parameter that adjusts the balance between the LQ performance and the actuation rate. While $\mu^{u}_{k}$ takes values in a continuous set, $\mu^{\delta}_{k}$ takes values in the binary set $\{0,1\}$. Consequently, Problem (14) is an optimal control problem with mixed discrete and continuous variables, which is generally difficult to solve directly. To effectively address this problem, we employ the rollout approach as an approximation technique.
III Proposed Rollout-Based algorithm
This section presents algorithm design using a rollout approach. In Section III-A, we give preliminaries on the rollout method. Section III-B derives the optimal periodic policy that is used in the algorithm design, and section III-C describes the algorithm design.
III-A Preliminaries for the Rollout Algorithm
The rollout algorithm is a sequential optimization technique in the context of dynamic programming [16, 17]. In this framework, the optimal value function of the original problem is approximated using a given policy, referred to as the base policy, and a suboptimal policy is obtained through an iterative procedure.
Since directly applying the rollout idea to the average-cost minimization problem (14) is analytically challenging, we instead consider the corresponding infinite-horizon discounted-cost problem, which serves as a tractable surrogate for constructing the rollout-based algorithm. Formally, we formulate the discounted-cost minimization problem as
$\underset{\displaystyle\mu^{u},\mu^{\delta}}{\mathrm{min}}\quad J^{d}(\mu^{u},\mu^{\delta})=J^{d}_{c}(\mu^{u},\mu^{\delta})+\theta J^{d}_{r}(\mu^{\delta})\hfil\hfil\hfil\hfil$ (15)
[…]
where
$J^{d}_{c}(\mu^{u},\mu^{\delta})$ $=\mathbb{E}\left[\sum_{k=0}^{\infty}\alpha^{k}\left(x_{k}^{\top}Qx_{k}+u_{k}^{\top}Ru_{k}\right)\right],$ (16)
$J^{d}_{r}(\mu^{\delta})$ $=\mathbb{E}\left[\sum_{k=0}^{\infty}\alpha^{k}\delta_{k}\right],$ (17)
and $\alpha\in(0,1)$ is the discount factor. By applying the dynamic programming principle, we obtain the following Bellman equation:
$V^{d}_{k}(\mathcal{I}_{k})=\min_{\begin{subarray}{c}u_{k},\dots,u_{k+h-1},\\ \delta_{k},\dots,\delta_{k+h-1}\end{subarray}}$ $\mathbb{E}\left[\sum_{i=k}^{k+h-1}\alpha^{i-k}g(x_{i},u_{i},\delta_{i})\right$.
$\quad\left.+\alpha^{h}V^{d}_{k+h}(\mathcal{I}_{k+h})\,\middle|\,\mathcal{I}_{k}\right],$ (18)
where $V^{d}_{k}$ is the optimal value function and $g(x,u,\delta)\coloneqq x^{\top}Qx+u^{\top}Ru+\theta\delta$.
Based on the rollout method, we replace $V^{d}_{k+h}$ in (18) by $\tilde{V}^{d}_{k+h}$, which is a value function associated with a fixed base policy. Since $\tilde{V}^{d}_{k+h}$ is independent of the optimization variables, the problem reduces to the following $h$-step lookahead minimization:
$\hat{V}^{d}_{k}(\mathcal{I}_{k})=\min_{\begin{subarray}{c}u_{k},\dots,u_{k+h-1},\\ \delta_{k},\dots,\delta_{k+h-1}\end{subarray}}$ $\mathbb{E}\left[\sum_{i=k}^{k+h-1}\alpha^{i-k}g(x_{i},u_{i},\delta_{i})\right$.
$\quad\left.+\alpha^{h}\tilde{V}^{d}_{k+h}(\mathcal{I}_{k+h})\,\middle|\,\mathcal{I}_{k}\right],$ (19)
At each time $k=\ell h$, the triggering policies over the $h$ steps are obtained by solving (19) with a fixed control policy in a receding-horizon manner. This yields a sequence of triggering policies $\mu^{\delta}_{k},\dots,\mu^{\delta}_{k+h-1}$, and the corresponding control policies $\mu^{u}_{k},\dots,\mu^{u}_{k+h-1}$ are then computed online. The procedure is repeated every $h$ steps.
In this study, we adopt the periodic policy, in which the control input is applied only at fixed time instants, as the base policy. Since the objective of this paper is to balance the trade-off between control performance and input sparsity, employing a base policy that yields intermittent actuation provides a consistent choice for approximating the value function. In what follows, we derive the optimal periodic policy for the discounted-cost minimization problem (15) and characterize the corresponding value function. For analytical purposes, we also derive the optimal periodic policy for Problem (14).
III-B Optimal Periodic Policy and The Corresponding Cost
Let $p\in\mathbb{N}$ denote the control period of the periodic policy. In this case, the associated triggering policy is given by
$\mu^{\delta}_{k}(\mathcal{I}_{k})=\begin{cases}1&\text{if }k\equiv 0\pmod{p},\\ 0&\text{otherwise},\end{cases}$ (20)
meaning that $u_{k}$ is applied every $p$ steps and $u_{k}=0$ otherwise. Hereafter, we denote the sequences of periodic control and triggering policies with period $p$ by $\mu^{u,\text{per}}$ and $\mu^{\delta,\text{per}}$, respectively.
For a nonnegative integer $\ell\in\mathbb{Z}_{\geq 0}$, we obtain the lifted system describing the state evolution from $k=\ell p$ to $k=(\ell+1)p$ as
$x_{(\ell+1)p}=A^{p}x_{\ell p}+B_{p}u_{\ell p}+D_{p}\bar{w}_{\ell}^{(p)},$ (21)
where
$B_{p}$ $\coloneqq A^{p-1}B,\quad D_{p}\coloneqq\begin{bmatrix}A^{p-1}&A^{p-2}&\cdots&I_{n_{x}}\end{bmatrix}$, (22)
$\bar{w}_{\ell}^{(p)}$ $\coloneqq\begin{bmatrix}w_{\ell p}^{\top}&w_{\ell p+1}^{\top}&\cdots&w_{(\ell+1)p-1}^{\top}\end{bmatrix}^{\top}$. (23)
The random vector $\bar{w}_{\ell}^{(p)}$ follows an independent Gaussian distribution with mean zero, whose covariance matrix $\Omega_{w}^{(p)}$ is given by
$\Omega_{w}^{(p)}=I_{p}\otimes\Omega_{w}.$ (24)
To reformulate the discounted-cost problem (15) and the average-cost problem (14) based on the lifted representation (21), we introduce the matrices $Q_{p}$, $S_{p}$, and $R_{p}$ as
$Q_{p}$ $\coloneqq\sum_{i=0}^{p-1}(A^{i})^{\top}QA^{i},\quad S_{p}\coloneqq\sum_{i=1}^{p-1}(A^{i})^{\top}QA^{i-1}B,$ (25)
$R_{p}$ $\coloneqq R+\sum_{i=1}^{p-1}B^{\top}(A^{i-1})^{\top}QA^{i-1}B.$ (26)
We also define the function $c_{p}\colon\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\to\mathbb{R}$ as $c_{p}(x,u)\coloneqq x^{\top}Q_{p}x+2x^{\top}S_{p}u+u^{\top}R_{p}u$. Then, using (21), $J^{d}_{c}$ and $J^{a}_{c}$ can be expressed as
$J^{d}_{c}(\mu^{u},\mu^{\delta,\text{per}})$ $=\mathbb{E}\left[\sum_{\ell=0}^{\infty}\alpha^{\ell p}c_{p}(x_{\ell p},u_{\ell p})\right]+d^{d}(p),$ (27)
$J^{a}_{c}(\mu^{u},\mu^{\delta,\text{per}})$ $=\limsup_{K\to\infty}\frac{1}{pK}\mathbb{E}\left[\sum_{\ell=0}^{K-1}c_{p}(x_{\ell p},u_{\ell p})\right]$
$\quad+\frac{1}{p}d^{a}(p),$ (28)
where $d^{d}(p)$ and $d^{a}(p)$ are given by
$d^{d}(p)$ $\coloneqq\frac{1}{1-\alpha^{p}}\sum_{i=1}^{p-1}\alpha^{i}\sum_{j=0}^{i-1}\operatorname{tr}\left(QA^{j}\Omega_{w}(A^{j})^{\top}\right),$ (29)
$d^{a}(p)$ $\coloneqq\sum_{i=1}^{p-1}\sum_{j=0}^{i-1}\operatorname{tr}\left(QA^{j}\Omega_{w}(A^{j})^{\top}\right).$ (30)
Since $\mu^{\delta,\text{per}}$ is fixed independently of $\mu^{u}$, the terms $J^{d}_{r}$ and $J^{a}_{r}$ can be computed explicitly as
$J^{d}_{r}(\mu^{\delta,\text{per}})=\frac{1}{1-\alpha^{p}},\quad J^{a}_{r}(\mu^{\delta,\text{per}})=\frac{1}{p}.$ (31)
Therefore, once $\mu^{\delta,\text{per}}$ is given as the triggering policy, Problem (15) reduces to finding $\mu^{u}$ that minimizes $J^{d}_{c}$. Similarly, for the average-cost problem (14), it suffices to find $\mu^{u}$ that minimizes $J^{a}_{c}$.
Suppose the following assumption holds.
Assumption 1
Any eigenvalues of $A$, denoted by $\lambda_{1}$ and $\lambda_{2}$, satisfy
$\lambda_{1}\neq\lambda_{2}\exp\left(\frac{2\pi q\sqrt{-1}}{p}\right),\quad\forall q\in\mathbb{Z}.$ (32)
This assumption ensures that $(A^{p},B_{p})$ is stabilizable when $(A,B)$ is controllable [20]. Then, the control policy that minimizes $J^{d}_{c}$ under control period $p$ can be obtained using standard optimal control techniques [21]:
$\mu^{u,\text{per}}_{k}=\begin{cases}\tilde{F}_{p}\hat{x}_{k}&\text{if }k\equiv 0\pmod{p},\\ 0&\text{otherwise},\end{cases}$ (33)
where
$\tilde{F}_{p}=-\left(\alpha^{p}B_{p}^{\top}\tilde{P}_{p}B_{p}+R_{p}\right)^{-1}\left(\alpha^{p}B_{p}^{\top}\tilde{P}_{p}A^{p}+S_{p}^{\top}\right),$ (34)
and $\tilde{P}_{p}$ is the positive-definite solution of the following algebraic Riccati equation [^2] If $(A,Q^{1/2})$ is observable, then $(A^{p},Q_{p}^{1/2})$ is also observable, which ensures the existence of a positive-definite solution (see Appendix -B).:
$\tilde{P}_{p}$ $=Q_{p}+\alpha^{p}(A^{p})^{\top}\tilde{P}_{p}A^{p}-M_{p},$ (35)
$M_{p}$ $\coloneqq\tilde{F}_{p}^{\top}\left(\alpha^{p}B_{p}^{\top}\tilde{P}_{p}B_{p}+R_{p}\right)\tilde{F}_{p}.$ (36)
The discounted cost associated with the periodic policy $(\mu^{u,\text{per}},\mu^{\delta,\text{per}})$ is given by
$J^{d}(\mu^{u,\text{per}},\mu^{\delta,\text{per}})=\mathbb{E}\left[x_{0}^{\top}\tilde{P}_{p}x_{0}\,\middle|\,\mathcal{I}_{0}\right]$
$\quad+\frac{\alpha^{p}}{1-\alpha^{p}}\operatorname{tr}\left(\tilde{P}_{p}D_{p}\Omega_{w}^{(p)}D_{p}^{\top}\right)+\sum_{\ell=0}^{\infty}\alpha^{\ell p}\operatorname{tr}\left(M_{p}\Sigma_{\ell p}\right)$
$\quad+d^{d}(p)+\frac{\theta}{1-\alpha^{p}}.$ (37)
Furthermore, the periodic control law that minimizes $J^{a}$ under control period $p$ is the same as (33), with $\alpha=1$ substituted into (34) and (35), and the associated average cost is calculated as follows:
$J^{a}(\mu^{u,\text{per}},\mu^{\delta,\text{per}})$ $=\frac{1}{p}\Big[\operatorname{tr}\left(\tilde{P}_{p}D_{p}\Omega_{w}^{(p)}D_{p}^{\top}\right)+\operatorname{tr}\left(M_{p}\Sigma\right)$ (38)
$\quad+d^{a}(p)\Big]+\frac{\theta}{p}.$ (39)
In the following section, we investigate the rollout-based algorithm for Problem (15) using the periodic policy $(\mu^{u,\text{per}},\mu^{\delta,\text{per}})$ as the base policy, and relate the results to Problem (14).
III-C Algorithm Design
Before proceeding to the algorithm design, we introduce some notation. Let $\mathcal{T}$ denote the set of sequences of triggering variables in which the first $h$ elements of $\{\delta_{k}\}_{k\in\mathbb{Z}_{\geq 0}}$ are free parameters, while the remaining elements follow the periodic triggering policy $\mu^{\delta,\text{per}}$. Since $\mathcal{T}$ contains $2^{h}$ possible sequences, there exists a bijective correspondence between $\mathcal{T}$ and the index set $\mathcal{M}\coloneqq\{1,2,\dots,2^{h}\}$. Hence, for each $m\in\mathcal{M}$, an arbitrary sequence in $\mathcal{T}$ can be represented as $\{\delta^{(m)}_{k}\}_{k\in\mathbb{Z}_{\geq 0}}$. With this notation, we characterize the triggering sequence $\{\delta^{(m)}_{k}\}_{k\in\mathbb{Z}_{\geq 0}}\in\mathcal{T},\ m\in\mathcal{M}$ as
$\delta^{(m)}_{k}$ $=\rho^{(m)}_{k},\quad k\in\{0,1,\dots,h-1\},$ (40)
$\delta^{(m)}_{k}$ $=\begin{cases}1&\text{if }k\equiv 0\pmod{p},\\ 0&\text{otherwise},\end{cases}\quad k\geq h,$ (41)
where $\rho^{(m)}_{k}\in\{0,1\}$ for all $k\in\mathbb{Z}_{\geq 0}$. In particular, for $m=1$, we define
$\rho^{(1)}_{k}=\begin{cases}1&\text{if }k\equiv 0\pmod{p},\\ 0&\text{otherwise},\end{cases}\quad k\in\{0,1,\dots,h-1\}.$ (42)
We are now ready to construct the algorithm for Problem (15) with the base policy $(\mu^{u,\text{per}},\mu^{\delta,\text{per}})$. Let $h$ be chosen so that $h/p\in\mathbb{N}$. From (37), the value function under the periodic policy $(\mu^{u,\text{per}},\mu^{\delta,\text{per}})$ for the discounted-cost problem (15) is obtained as
$V^{d,\text{per}}_{\ell p}(\mathcal{I}_{\ell p})=\mathbb{E}\left[x_{\ell p}^{\top}\tilde{P}_{p}x_{\ell p}\,\middle|\,\mathcal{I}_{\ell p}\right]$
$\quad+\frac{\alpha^{p}}{1-\alpha^{p}}\operatorname{tr}\left(\tilde{P}_{p}D_{p}\Omega_{w}^{(p)}D_{p}^{\top}\right)+\sum_{j=\ell}^{\infty}\alpha^{(j-\ell)p}\operatorname{tr}\left(M_{p}\Sigma_{jp}\right)$
$\quad+d^{d}(p)+\frac{\theta}{1-\alpha^{p}},$ (43)
for $k=\ell p$. Since $h/p\in\mathbb{N}$, we have $k+h\equiv 0\pmod{p}$ when $k=\ell p$. Substituting $V^{d,\text{per}}_{k+h}$ for $V^{d}_{k+h}$ in (19) and rewriting $V^{d}_{k}$ as $\hat{V}^{d}_{k}$, we obtain the following equation:
$\hat{V}^{d}_{k}(\mathcal{I}_{k})=\min_{\begin{subarray}{c}u_{k},\dots,u_{k+h-1},\\ \delta_{k},\dots,\delta_{k+h-1}\end{subarray}}\mathbb{E}\left[\sum_{i=k}^{k+h-1}\alpha^{i-k}g(x_{i},u_{i},\delta_{i})\right$.
$\quad\left.+\alpha^{h}\mathbb{E}\left[x_{k+h}^{\top}\tilde{P}_{p}x_{k+h}\,\middle|\,\mathcal{I}_{k+h}\right]\,\middle|\,\mathcal{I}_{k}\right]+\alpha^{h}\eta^{d}_{k},$ (44)
where
$\eta^{d}_{k}\coloneqq\frac{\alpha^{p}}{1-\alpha^{p}}\operatorname{tr}\left(\tilde{P}_{p}D_{p}\Omega_{w}^{(p)}D_{p}^{\top}\right)$
$+\sum_{j=(k+h)/p}^{\infty}\alpha^{(j-(k+h)/p)p}\operatorname{tr}\left(M_{p}\Sigma_{jp}\right)+d^{d}(p)+\frac{\theta}{1-\alpha^{p}}.$ (45)
By the tower property of conditional expectation [22, Th. 5.1.6], we have
$\mathbb{E}\left[\mathbb{E}\left[x_{k+h}^{\top}\tilde{P}_{p}x_{k+h}\,\middle|\,\mathcal{I}_{k+h}\right]\,\middle|\,\mathcal{I}_{k}\right]=\mathbb{E}\left[x_{k+h}^{\top}\tilde{P}_{p}x_{k+h}\,\middle|\,\mathcal{I}_{k}\right].$ (46)
Since $\eta^{d}_{k}$ does not depend on $\mu^{u}$ and $\mu^{\delta}$, it suffices to minimize the following cost to obtain the policy minimizing $\hat{V}^{d}_{k}$ in (44):
$\mathbb{E}\left[\sum_{i=k}^{k+h-1}\alpha^{i-k}g(x_{i},u_{i},\delta_{i})+\alpha^{h}x_{k+h}^{\top}\tilde{P}_{p}x_{k+h}\,\middle|\,\mathcal{I}_{k}\right].$ (47)
Note that if $m\in\mathcal{M}$ is fixed, this cost (47) is a standard finite-horizon discounted cost conditioned on $\mathcal{I}_{k}$. Therefore, for a fixed $m\in\mathcal{M}$, we can derive the control input minimizing (47) as
$u_{k}=\rho^{(m)}_{s}F^{(m)}_{s}\hat{x}_{k},\quad s\in\{0,1,\dots,h-1\},$ (48)
where for any $s\in\{0,1,\dots,h-1\}$ and $m\in\mathcal{M}$,
$F^{(m)}_{s}=-\alpha\big(\alpha B^{\top}P^{(m)}_{s+1}B+R\big)^{-1}B^{\top}P^{(m)}_{s+1}A,$ (49)
and $P^{(m)}_{s}$ is given recursively by
$P^{(m)}_{h}$ $=\tilde{P}_{p},$ (50)
$P^{(m)}_{s}$ $=Q+\alpha A^{\top}P^{(m)}_{s+1}A$
$\quad-\alpha^{2}A^{\top}P^{(m)}_{s+1}B\left(\alpha B^{\top}P^{(m)}_{s+1}B+R\right)^{-1}B^{\top}P^{(m)}_{s+1}A,$
$\quad s\in\{0,1,\dots,h-1\}.$ (51)
Hence, for a given $m\in\mathcal{M}$, $\hat{V}^{d}_{k}$ can be expressed as
$\hat{V}^{d}_{k}(\mathcal{I}_{k})=\hat{x}_{k}^{\top}P^{(m)}_{0}\hat{x}_{k}+\operatorname{tr}\left(P^{(m)}_{0}\Sigma_{k}\right)+\beta^{(m)}_{k}+\gamma^{(m)},$ (52)
where for each $m\in\mathcal{M}$,
$\beta^{(m)}_{k}$ $\coloneqq\sum_{\tau=0}^{h-1}\alpha^{\tau}\left(\operatorname{tr}\left(P^{(m)}_{\tau+1}\Omega_{w}\right)+\operatorname{tr}\left(M^{(m)}_{\tau}\Sigma_{k+\tau}\right)\right),$ (53)
$M^{(m)}_{\tau}$ $\coloneqq(F^{(m)}_{\tau})^{\top}\left(\alpha B^{\top}P^{(m)}_{\tau+1}B+R\right)F^{(m)}_{\tau}$, (54)
$\gamma^{(m)}$ $\coloneqq\theta\sum_{\tau=0}^{h-1}\alpha^{\tau}\rho^{(m)}_{\tau}$. (55)
Therefore, by selecting $m\in\mathcal{M}$ that minimizes $\hat{V}^{d}_{k}$ in (52) and computing the corresponding control input using (48), we can construct the policies over the interval [k,k+h1][k,k+h-1].
The preceding discussion has focused on the discounted-cost problem (15) with $\alpha\in(0,1)$. We now construct the algorithm for the average-cost problem (14) by taking the limit as $\alpha\uparrow 1$ in (44). Although $\eta^{d}_{k}$ diverges in this limit, it does not affect the optimization procedure since it is independent of the optimization variables. Furthermore, since $(A^{p},B_{p})$ is stabilizable, $(A^{p},Q_{p}^{1/2})$ is observable, and $R_{p}$ is positive definite, a stabilizing solution $\tilde{P}_{p}$ exists at $\alpha=1$, and $\tilde{P}_{p}$ is continuous on $\alpha\in(0,1]$. Consequently, $P^{(m)}_{k}$, $F^{(m)}_{k}$, and $M^{(m)}_{k}$ are also continuous on $\alpha\in(0,1]$, and hence (48) and (52) remain valid at $\alpha=1$.
Summarizing the above discussion, the proposed algorithm for Problem (14) is presented as follows.
Algorithm 1
  1. (i)
    At time $k=\ell h,\ \ell\in\mathbb{Z}_{\geq 0}$, compute
    $\varphi(\mathcal{I}_{\ell h})\coloneqq\underset{m\in\mathcal{M}}{\operatorname{arg\,min}}\$$\hat{x}_{\ell h}^{\top}P^{(m)}_{0}\hat{x}_{\ell h}+\operatorname{tr}\left(P^{(m)}_{0}\Sigma_{\ell h}\right)$
    $\quad+\beta^{(m)}_{\ell h}+\gamma^{(m)}.$ (56)
  2. (ii)
    Determine the triggering variables over the interval $\ell h,(\ell+1)h-1]$ by
    $\delta_{k}=\rho^{(\varphi(\mathcal{I}_{\ell h}))}_{k-\ell h}$, (57)
    and compute the corresponding control input as
    $u_{k}=\rho^{(\varphi(\mathcal{I}_{\ell h}))}_{k-\ell h}F^{(\varphi(\mathcal{I}_{\ell h}))}_{k-\ell h}\hat{x}_{k}.$ (58)
  3. (iii)
    Repeat steps (i) and (ii) every $h$ steps.
Therefore, the control policy $\mu^{u}_{k}$ and the triggering policy $\mu^{\delta}_{k}$ are determined by Algorithm 1 as
$(\mu^{u}_{k}(\mathcal{I}_{k}),\mu^{\delta}_{k}(\mathcal{I}_{k}))$ $=\left(\rho^{(\varphi(\mathcal{I}_{\ell h}))}_{k-\ell h}F^{(\varphi(\mathcal{I}_{\ell h}))}_{k-\ell h}\hat{x}_{k},\rho^{(\varphi(\mathcal{I}_{\ell h}))}_{k-\ell h}\right)$, (59)
$\ell h\leq k<(\ell+1)h,\quad\ell\in\mathbb{Z}_{\geq 0}.$ (60)
IV Performance and Stability Guarantees of the Proposed Algorithm
This section establishes theoretical performance guarantees for Algorithm 1 in Problem (14), in comparison with periodic control. We also provide a stability result for the closed-loop system under the policy obtained by Algorithm 1.
We impose the following assumptions for the analysis.
Assumption 2
  1. (i)
    $\operatorname{rank}(C)=\operatorname{rank}(C^{\top})=n_{x}$.
  2. (ii)
    $Q\succ 0$.
  3. (iii)
    $\Omega_{x_{0}}$ is a solution of the following Riccati equation:
    $\Omega_{x_{0}}$ $=\Omega_{w}+A\Omega_{x_{0}}A^{\top}$ (61)
    $-A\Omega_{x_{0}}C^{\top}\left(C\Omega_{x_{0}}C^{\top}+\Omega_{v}\right)^{-1}C\Omega_{x_{0}}A^{\top}.$ (62)
Assumption 2(i) ensures that the steady-state Kalman gain $G$ has full row rank, since the covariance matrix $\tilde{\Sigma}$ is positive definite. Assumption 2(ii) guarantees that the minimum eigenvalue of $Q$ is strictly positive [23, Lemma 10.4.1]. When Assumption 2(iii) holds, $\Omega_{x_{0}}$ is positive definite since $(A,\Omega_{w}^{1/2})$ is controllable. Furthermore, this assumption implies $G_{k}=G$ for all $k$, i.e., the Kalman filter is stationary, which is the case of interest in this study.
Let $(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}})$ denote the sequence of the policies obtained by Algorithm 1. The following theorem provides a theoretical performance bound.
Theorem 1
Suppose that Assumptions 1 and 2 hold. Then, the following inequality holds:
$J^{a}(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}})\leq J^{a}(\mu^{u,\text{per}},\mu^{\delta,\text{per}})+\frac{1}{h}.$ (63)
Proof:
See Appendix -C. ∎
The additive term $1/h$ in (63) appears due to technical reasons. When $h$ is sufficiently large, this term becomes negligible. In such cases, Algorithm 1 is executed only at sparse time instants, with most actuation timings and feedback gains being computed in an open-loop fashion. Consequently, the control policy $\mu^{u,\text{ro}}_{k}$ and the triggering policy $\mu^{\delta,\text{ro}}_{k}$ are determined primarily by statistical information, such as $\Omega_{w}$, $\Omega_{v}$, and $\Sigma$, rather than by actual disturbances. This resembles the optimal periodic control, where the periodic policy is determined from stationary statistics via minimizing (39). In such situations, (63) shows that Algorithm 1 guarantees performance no worse than the optimal periodic policy.
For a given $m\in\mathcal{M}$, we define $V_{m}\colon\mathbb{R}^{n_{x}}\to[0,\infty)$ by
$V_{m}(x)\coloneqq x^{\top}P_{0}^{(m)}x,\quad\forall x\in\mathbb{R}^{n_{x}}.$ (64)
For notational convenience, we introduce $\chi_{\ell}\coloneqq x_{\ell h}$ and $\hat{\chi}_{\ell}\coloneqq\hat{x}_{\ell h}$ for $\ell\in\mathbb{Z}_{\geq 0}$. The following lemmas are used to establish Theorem 1.
Lemma 1
If Assumptions 1 and 2 hold, then it holds that
$\limsup_{L\to\infty}\mathbb{E}\left[V_{1}(\chi_{L})\,\middle|\,\hat{\chi}_{0}\right]<\infty.$ (65)
Proof:
See Appendix -D. ∎
Lemma 2
Under Assumptions 1 and 2, the stochastic process $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ induced by $(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}})$ is a Markov chain.
Proof:
See Appendix -E. ∎
We define $f_{m}\colon\mathbb{R}^{n_{x}}\to[0,\infty)$ as
$f_{m}(x)$ $\coloneqq x^{\top}\left(P_{0}^{(1)}-P_{0}^{(m)}\right)x+\operatorname{tr}\left(\left(P_{0}^{(1)}-P_{0}^{(m)}\right)\Sigma\right)$
$\quad+\beta^{(1)}-\beta^{(m)}+\gamma^{(1)}-\gamma^{(m)}.$ (66)
In (66), we write $\beta^{(m)}$ instead of $\beta^{(m)}_{k}$, since $\beta^{(m)}_{k}$ is independent of the time index $k$ under Assumption 2. By construction of Algorithm 1, we have $f_{m}(x)\geq 0$ for all $x\in\mathbb{R}^{n_{x}}$. Then, it follows that the Markov chain $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ is ergodic, i.e., positive Harris recurrent [24, Ch. 14].
Theorem 2
Define the mapping $\tilde{f}_{m}\coloneqq f_{m}+1$. Let Assumptions 1 and 2 hold. Then, under the policy $(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}})$, the Markov chain $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ is $\tilde{f}_{m}$-ergodic.
Proof:
See Appendix -F. ∎
Furthermore, Lemma 1 implies that Algorithm 1 ensures mean-square stability of the closed-loop system, as stated below.
Theorem 3
Suppose that Assumptions 1 and 2 hold. Then, the closed-loop system under $(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}})$ is mean-square stable, i.e.,
$\sup_{k\in\mathbb{Z}_{\geq 0}}\mathbb{E}\left[\|x_{k}\|^{2}\right]<\infty.$ (67)
Proof:
See Appendix -G. ∎
V Illustrative Example
Consider a system consisting of two masses $m_{1}$ and $m_{2}$ connected by a spring with spring constant $\kappa_{s}$ [18]. We set $m_{1}=m_{2}=1$ and $\kappa_{s}=2\pi^{2}$. The control input and the process noise affect only mass $m_{1}$. The model of this system is described by the following stochastic differential equations:
$dx(t)$ $=\left(A_{C}x(t)+B_{C}u(t)\right)dt+D_{C}dw(t),$ (68)
$dy(t)$ $=C_{C}x(t)dt+dv(t),$ (69)
where
$A_{C}$ $=\begin{bmatrix}0&0&1&0\\ 0&0&0&1\\ -\frac{\kappa_{s}}{m_{1}}&\frac{\kappa_{s}}{m_{1}}&0&0\\ \frac{\kappa_{s}}{m_{2}}&-\frac{\kappa_{s}}{m_{2}}&0&0\end{bmatrix},\quad B_{C}=\begin{bmatrix}0\\ 0\\ \frac{1}{m_{1}}\\ 0\end{bmatrix}$, (70)
$D_{C}$ $=\begin{bmatrix}0\\ 0\\ 0.4\\ 0\end{bmatrix},\quad C_{C}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\end{bmatrix}$, (71)
and $x(t)=[x^{1}(t)\ x^{2}(t)\ \dot{x}^{1}(t)\ \dot{x}^{2}(t)]^{\top}\in\mathbb{R}^{4}$ denotes the state vector, where $x^{1}(t)$ and $x^{2}(t)$ are the displacements of masses $m_{1}$ and $m_{2}$, respectively. We assume that $w(t)$ and $v(t)$ are independent Wiener processes with zero mean, whose incremental covariances are $I_{n_{x}}dt$ and $10^{-5}I_{n_{x}}dt$, respectively. By discretizing (68) and (69) with the sampling period $t_{s}=0.1$, we obtain the discrete-time system (2) and (3). The discretized disturbances $w_{k}$ and $v_{k}$ are i.i.d. Gaussian random vectors with zero mean and covariance matrices
$\Omega_{w}=\int_{0}^{t_{s}}e^{A_{C}\tau}D_{C}D_{C}^{\top}e^{A_{C}^{\top}\tau}\,d\tau,\quad\Omega_{v}=\frac{1}{t_{s}}\,E.$ (72)
For the discretized system, we choose $\bar{x}_{0}=[1\ -1\ 0\ 0]^{\top}$ and let $\Omega_{x_{0}}$ satisfy (62). We set the control horizon to 600, and $Q$ and $R$ are given by
$Q$ $=\begin{bmatrix}0.1336&-0.0936&-0.0327&0.0347\\ -0.0936&0.1336&0.0347&-0.0327\\ -0.0327&0.0347&0.0377&0.0024\\ 0.0347&-0.0327&0.0024&0.0377\end{bmatrix}$, (73)
$R$ $=0.1.$ (74)
We evaluate three methods for Problem (14). The first one is Algorithm 1 presented in Section III-A, with $h=6$. At $k=6\ell$ for $\ell\in\mathbb{Z}_{\geq 0}$, the triggering variables and the corresponding control inputs are computed according to (60). The second method is the periodic control described in Section III-A, with $p\in\{1,2,3,6\}$. The third method is an $\ell_{1}$-relaxation approach, where $\delta_{k}$ in (13) is replaced by $\|u_{k}\|_{\ell_{2}/\ell_{1}}$, known as the $\ell_{2}/\ell_{1}$ norm [25]. To ensure robustness against disturbances, we employ an MPC scheme with a prediction horizon of 30 and solve the $\ell_{1}$-relaxed problem at each time step using CVX [26]. This method is hereafter referred to as $\ell_{1}$-relax.+MPC. We conduct Monte Carlo simulations for each method using different values of $\theta\in\{0.02,0.04,\dots,0.40\}$, and compare the means of the control cost and the actuation rate across trials for each $\theta$.
Fig. 2 shows the average control cost versus the average actuation rate, and Fig. 3 presents each cost as a function of $\theta$. In Fig. 3, error bars indicate one standard deviation around the mean over 50 trials. For the periodic control, $p\in\{1,2,3,6\}$ is chosen so that the average cost associated with the periodic policy (39) is minimized for each $\theta$. Compared with periodic control, the proposed algorithm achieves lower control costs for all $\theta$, while maintaining nearly the same average actuation rate. Although $\ell_{1}$-relax.+MPC yields better control performance than the proposed method, its average actuation rate is relatively high. Thus, our method provides a superior trade-off compared with the other two approaches, highlighting the effectiveness of the rollout-based algorithm we propose.
Figure 2: Plot of the average control cost versus the average actuation rate.
Figure 3: Average control cost (top) and average actuation rate (bottom) as functions of $\theta$, with error bars indicating the standard deviation from the mean.
VI Conclusion
We have proposed a co-design method of control and triggering policies that balances the trade-off between control performance and input sparsity. To this end, we formulated the optimal control problem with actuation-rate regularization and developed a rollout-based algorithm using a receding horizon approach. The proposed algorithm provides a theoretical performance guarantee and ensures mean-square stability of the closed-loop system. Its effectiveness has been demonstrated through a numerical simulation. Future work includes establishing performance guarantees for the discounted-cost problem and deriving strict performance improvement results over the periodic control strategy.
This appendix provides some preliminaries as well as the detailed proofs of the analytical results established in this paper.
-A Preliminaries for the Proofs
When Assumption 2 holds, we can write $\varphi$ as
$\varphi(\mathcal{I}_{\ell h})=\underset{m\in\mathcal{M}}{\operatorname{arg\,min}}\$$\hat{x}_{\ell h}^{\top}P^{(m)}_{0}\hat{x}_{\ell h}+\operatorname{tr}\left(P^{(m)}_{0}\Sigma\right)$
$\quad+\beta^{(m)}+\gamma^{(m)}.$ (75)
[…]
Next, we derive the dynamics of $\hat{\chi}_{\ell}$ induced by the policies $(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}})$. From (7), we have
$\hat{x}_{k+1}=(A+\rho_{k-\ell h}^{(\varphi(\hat{\chi}_{\ell}))}BF_{k-\ell h}^{(\varphi(\hat{\chi}_{\ell}))})\hat{x}_{k}$ (76)
$\quad+G\left(y_{k+1}-C(A+\rho_{k-\ell h}^{(\varphi(\hat{\chi}_{\ell}))}BF_{k-\ell h}^{(\varphi(\hat{\chi}_{\ell}))})\hat{x}_{k}\right).$ (77)
For $i\in\{0,1,\dots,h-1\}$ and $m\in\mathcal{M}$, we define
$\Theta_{i}^{(m)}$ $\coloneqq A+\rho_{i}^{(m)}BF_{i}^{(m)},$ (78)
$\omega_{k}$ $\coloneqq G\left(CAe_{k}+Cw_{k}+v_{k+1}\right).$ (79)
Then, Eq. (77) can be rewritten as
$\hat{x}_{k+1}=\Theta_{k-lh}^{(\varphi(\hat{\chi}_{\ell}))}\hat{x}_{k}+\omega_{k}.$ (80)
By iterating this equation from $k=\ell h$ to $k=(\ell+1)h-1$, we obtain
$\hat{\chi}_{\ell+1}=\Phi_{\varphi(\hat{\chi}_{\ell})}\hat{\chi}_{\ell}+\Gamma_{\varphi(\hat{\chi}_{\ell})}\bar{\omega}_{\ell},\quad\forall\ell\in\mathbb{Z}_{\geq 0},$ (81)
where
$\Phi_{m}$ $\coloneqq\prod_{s=h-1}^{0}\Theta_{s}^{(m)},$ (82)
$\Gamma_{m}$ $\coloneqq\begin{bmatrix}\prod_{s=h-1}^{1}\Theta_{s}^{(m)}&\prod_{s=h-1}^{2}\Theta_{s}^{(m)}&\dots&\Theta_{h-1}^{(m)}&I_{n_{x}}\end{bmatrix}$, (83)
$\bar{\omega}_{\ell}$ $\coloneqq\begin{bmatrix}\omega_{\ell h}^{\top}&\omega_{\ell h+1}^{\top}&\dots&\omega_{(\ell+1)h-1}^{\top}\end{bmatrix}^{\top}$. (84)
Finally, we rewrite the objective function $J^{a}$ under the policies $(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}})$. For a given $\mu^{\delta,\text{ro}}$, $J^{a}_{r}$ is finite, and $J^{a}_{c}$ converges to some finite value for any $(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}})$, which is the direct result from Theorem 1. Hence, we have
$J^{a}(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}})$ (85)
$=\limsup_{N\to\infty}\frac{1}{N}\mathbb{E}\left[\sum_{k=0}^{N-1}x_{k}^{\top}Qx_{k}+u_{k}^{\top}Ru_{k}+\theta\delta_{k}\right]$ (86)
$=\limsup_{L\to\infty}\frac{1}{hL}\mathbb{E}\Biggl[\sum_{\ell=0}^{L-1}\sum_{\tau=0}^{h-1}x_{\ell h+\tau}^{\top}Qx_{\ell h+\tau}+u_{\ell h+\tau}^{\top}Ru_{\ell h+\tau}$ (87)
$\qquad+\theta\delta_{\ell h+\tau}\Biggr]$ (88)
$\limsup_{L\to\infty}\frac{1}{hL}\mathbb{E}\left[\sum_{\ell=0}^{L-1}\bar{g}\left(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell}\right)\right],$ (89)
where
$\bar{g}(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell})$ (90)
[…]
and $\bar{w}_{\ell}\coloneqq[w_{\ell h}^{\top}\ w_{\ell h+1}^{\top}\ \dots w_{(\ell+1)h-1}^{\top}]^{\top}$.
-B Preliminary Lemmas
Lemma 3
If $(A,Q^{1/2})$ is observable, then for any $p\in\mathbb{N}$, the pair $(A^{p},Q_{p}^{1/2})$ is also observable.
Proof:
Let $\lambda_{p}\in\mathbb{C}$ be an eigenvalue of $A^{p}$ and $\xi_{p}\in\mathbb{C}^{n_{x}}\setminus\{0\}$ be a corresponding eigenvector, i.e., $A^{p}\xi_{p}=\lambda_{p}\xi_{p}$. Suppose that there exists such a vector $\xi_{p}$ satisfying $Q_{p}^{1/2}\xi_{p}=0$. From the definition of $Q_{p}$, it follows that
$0=\xi_{p}^{\ast}Q_{p}\xi_{p}=\sum_{i=0}^{p-1}(A^{i}\xi_{p})^{\ast}QA^{i}\xi_{p}=\sum_{i=0}^{p-1}\|Q^{1/2}A^{i}\xi_{p}\|^{2}.$ (92)
Hence, $Q^{1/2}A^{i}\xi_{p}=0$ for all $i=0,1,\dots,p-1$. Writing $s=qp+r$ with $q\in\mathbb{N}$ and $r\in\{0,1,\dots,p-1\}$, we obtain
$A^{s}\xi_{p}=A^{qp+r}\xi_{p}=\lambda_{p}^{q}A^{r}\xi_{p}.$ (93)
Since $Q^{1/2}A^{r}\xi_{p}=0$ for all $r=0,1,\dots,p-1$, Eq. (93) implies that $Q^{1/2}A^{s}\xi_{p}=0$ for all $s\in\mathbb{N}$. Consequently,
$\xi_{p}\in\{x\in\mathbb{C}^{n_{x}}\colon Q^{1/2}A^{s}x=0,\;\forall s\in\mathbb{N}\}.$ (94)
On the other hand, since $(A,Q^{1/2})$ is observable, the Cayley–Hamilton theorem ensures that
$\{x\in\mathbb{C}^{n_{x}}\colon Q^{1/2}A^{s}x=0,\;\forall s\in\mathbb{N}\}=\{0\}.$ (95)
This contradicts $\xi_{p}\neq 0$. Therefore, $Q_{p}^{1/2}\xi_{p}\neq 0$ holds for all eigenvectors $\xi_{p}$ of $A^{p}$, which completes the proof. ∎
Lemma 4
Suppose that $(A,Q^{1/2})$ is observable and that Assumption 1 holds for some $p\in\mathbb{N}$. Then, the matrix $\tilde{P}_{p}$ obtained by solving (35) with $\alpha=1$ is positive definite.
Proof:
Assumption 1 and Lemma 3 ensure that $\tilde{P}_{p}$ is positive semidefinite at $\alpha=1$ [21, Sec. 8.4]. Assume that there exists a nonzero vector $\xi\in\mathbb{R}^{n_{x}}$ such that $\xi^{\top}\tilde{P}_{p}\xi=0$. Then, we consider the following optimal control problem for a noise-free system with the initial condition $x_{0}=\xi$:
$\underset{\displaystyle\{u_{k}\}_{k\in\mathbb{Z}_{\geq 0}}}{\mathrm{min}}\quad\sum_{k=0}^{\infty}x_{k}^{\top}Q_{p}x_{k}+2x_{k}^{\top}S_{p}u_{k}+u_{k}^{\top}R_{p}u_{k}\hfil\hfil\hfil\hfil{}$ (96)
[…]
$x_{0}$ $=\xi.$
Since $(A^{p},B_{p})$ is stabilizable and $(A^{p},Q_{p}^{1/2})$ is observable by Lemma 3, the optimal objective value of (96) is $\xi^{\top}\tilde{P}_{p}\xi=0$. Furthermore, since $Q_{p}\succeq 0$ and $R_{p}\succ 0$, it follows that
$x_{k}^{\top}Q_{p}x_{k}+2x_{k}^{\top}S_{p}u_{k}+u_{k}^{\top}R_{p}u_{k}=0$ (97)
for all $k\in\mathbb{Z}_{\geq 0}$, which implies
$Q_{p}^{1/2}x_{k}=0,\quad u_{k}=0,\quad\forall k\in\mathbb{Z}_{\geq 0}.$ (98)
Since $u_{k}=0$ for all $k\in\mathbb{Z}_{\geq 0}$, it holds that $x_{k}=A^{pk}\xi$ for all $k\in\mathbb{Z}_{\geq 0}$. Consequently, $Q_{p}^{1/2}A^{pk}\xi=0$ for all $k\in\mathbb{Z}_{\geq 0}$, which contradicts the observability of $(A^{p},Q_{p}^{1/2})$ established in Lemma 3. Therefore, only $\xi=0$ satisfies $\xi^{\top}\tilde{P}_{p}\xi=0$, and thus $\tilde{P}_{p}$ is positive definite at $\alpha=1$. ∎
Lemma 5
Consider Algorithm 1 under Assumptions 1 and 2. Then, it holds that $P^{(1)}_{0}=\tilde{P}_{p}$.
Proof:
First, we can observe that $\hat{V}^{d}_{k}$ in (52) is obtained by applying the control policy $\mu^{u}_{\text{ro}}$ to the $h$-stage cost (47), starting from $k=\ell h$ for $\ell\in\mathbb{Z}_{\geq 0}$.
Define the following function:
$J_{h}(\mathcal{I}_{k})$ $\coloneqq\mathbb{E}\Biggl[x_{k+h}^{\top}\tilde{P}_{p}x_{k+h}$ (99)
$\quad+\sum_{i=k}^{k+h-1}\bigl(x_{i}^{\top}Qx_{i}+u_{i}^{\top}Ru_{i}+\theta\delta_{i}\bigr)\,\bigg|\,\mathcal{I}_{k}\Biggr].$ (100)
Then, we consider the following $h$-stage cost minimization problem:
$\underset{\displaystyle\{u_{i}\}_{i=k}^{k+h-1},\{\delta_{i}\}_{i=k}^{k+h-1}}{\mathrm{min}}\quad J_{h}(\mathcal{I}_{k})\hfil\hfil\hfil\hfil{}$ (101)
[…]
$y_{i}$ $=Cx_{i}+v_{i},\hfil\hfil$
$\delta_{i}$ $\in\{0,1\},\hfil\hfil$
$u_{i}$ $=0\quad\text{if }\delta_{i}=0.$
Let $H\coloneqq h/p\in\mathbb{N}$. If the periodic triggering policy with period $p$ is optimal for (101), then $J_{h}$ can be rewritten as
$J_{h}(\mathcal{I}_{k})$ $=\mathbb{E}\Biggl[x_{k+h}^{\top}\tilde{P}_{p}x_{k+h}+\sum_{j=0}^{H-1}\sum_{i=0}^{p-1}x_{k+jp+i}^{\top}Qx_{k+jp+i}$
$\qquad+u_{k+jp+i}^{\top}Ru_{k+jp+i}+\theta\delta_{k+jp+i}\,\Big|\,\mathcal{I}_{k}\Biggr]$ (102)
$=\mathbb{E}\Biggl[x_{k+h}^{\top}\tilde{P}_{p}x_{k+h}+\sum_{j=0}^{H-1}x_{k+jp}^{\top}Q_{p}x_{k+jp}$
$\qquad+2x_{k+jp}^{\top}S_{p}u_{k+jp}+u_{k+jp}^{\top}R_{p}u_{k+jp}\,\Big|\,\mathcal{I}_{k}\Biggr]$ (103)
$\quad+Hd^{a}(p)+H\theta.$ (104)
By applying the dynamic programming principle, we obtain the following recursion:
$P_{h}$ $=\tilde{P}_{p},$ (105)
$P_{t}$ $=Q_{p}+(A^{p})^{\top}P_{t+1}A^{p}-\left((A^{p})^{\top}P_{t+1}B_{p}+S_{p}\right)$ (106)
$\qquad\times\left(B_{p}^{\top}P_{t+1}B_{p}+R_{p}\right)^{-1}\left(B_{p}^{\top}P_{t+1}A^{p}+S_{p}^{\top}\right),$
$\quad t\in\{0,1,\dots,H-1\}.$ (107)
Comparing (107) with (35) under $\alpha=1$ shows that $P_{t}=\tilde{P}_{p}$ for all $t\in\{0,1,\dots,H-1\}$. Therefore, the optimal value of (101) corresponding to the periodic policy is given by
$J_{h}(\mathcal{I}_{k})=\hat{x}_{k}^{\top}\tilde{P}_{p}\hat{x}_{k}+\operatorname{tr}\left(\tilde{P}_{p}\Sigma\right)+\beta^{(1)}+\gamma^{(1)},$ (108)
where
$\beta^{(1)}$ $=H\Bigl(\operatorname{tr}\left(\tilde{P}_{p}D_{p}\Omega_{w}^{(p)}D_{p}^{\top}\right)+\operatorname{tr}\left(M_{p}\Sigma\right)+d^{a}(p)\Bigr),$ (109)
$\gamma^{(1)}$ $=H\theta.$ (110)
Comparing (52) and (108), we obtain $P^{(1)}_{0}=\tilde{P}_{p}$. ∎
Lemma 6
The function $V_{1}$ defined in (64) is coercive; that is, $V_{1}(x)\to\infty$ as $\|x\|\to\infty$.
Proof:
To establish the coerciveness of $V_{1}$, it is sufficient to show that the sublevel sets $\mathcal{L}_{r}(x)\coloneqq\{x\in\mathbb{R}^{n_{x}}\colon V_{1}(x)\leq r\}$ are precompact, i.e., their closures are compact, for each $r\in\mathbb{R}_{>0}$ [24, Sec. 9.4]. Since $P^{(1)}_{0}$ is positive definite by Lemmas 4 and 5, we have
$\mathcal{L}_{r}(x)\subseteq\left\{x\in\mathbb{R}^{n_{x}}\colon\|x\|^{2}\leq\frac{r}{\lambda_{\min}(P^{(1)}_{0})}\right\}$. (111)
The set on the right-hand side is bounded for every $r\in\mathbb{R}_{>0}$, and hence $\mathcal{L}_{r}(x)$ is also bounded. Moreover, the continuity of $V_{1}$ implies that $\mathcal{L}_{r}(x)=V_{1}^{-1}([0,r])$ is closed [27, Th. 4.8]. By the Heine–Borel theorem, $\mathcal{L}_{r}(x)$ is compact [27, Th. 2.41], and thus its closure is also compact for each $r\in\mathbb{R}_{>0}$ [27, Th. 2.27], which establishes the claim. ∎
-C Proof of Theorem 1
For the policies $(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}})$ obtained via Algorithm 1, we obtain, at iteration $\ell$,
$\mathbb{E}\Biggl[\sum_{i=0}^{h-1}x_{\ell h+i}^{\top}Qx_{\ell h+i}+u_{\ell h+i}^{\top}Ru_{\ell h+i}+\theta\delta_{\ell h+i}$ (112)
$\quad+x_{(\ell+1)h}\tilde{P}_{p}x_{(\ell+1)h}\,\Big|\,\mathcal{I}_{\ell h}\Biggr]$ (113)
[…]
$\quad+\theta\mu^{\delta,\text{ro}}_{\ell h+i}(\hat{\chi}_{\ell})+x_{(\ell+1)h}\tilde{P}_{p}x_{(\ell+1)h}\,\Big|\,\hat{\chi}_{\ell}\Biggr$] (115)
$=\mathbb{E}\left[V_{1}(\chi_{\ell+1})+\bar{g}\left(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell}\right)\,\middle|\,\hat{\chi}_{\ell}\right$] (116)
$=\hat{\chi}_{\ell}P^{(\varphi(\hat{\chi}_{\ell}))}_{0}\hat{\chi}_{\ell}+\operatorname{tr}\left(P^{(\varphi(\hat{\chi}_{\ell}))}_{0}\Sigma\right)+\beta^{(\varphi(\hat{\chi}_{\ell}))}+\gamma^{(\varphi(\hat{\chi}_{\ell}))}$. (117)
From (64) and the fact that
$\mathbb{E}\left[\chi_{\ell}^{\top}P^{(m)}_{0}\chi_{\ell}\,\middle|\,\hat{\chi}_{\ell}\right]=\hat{\chi}_{\ell}^{\top}P^{(m)}_{0}\hat{\chi}_{\ell}+\operatorname{tr}\left(P^{(m)}_{0}\Sigma\right),$ (118)
holds for all $m\in\mathcal{M}$ [28], it follows that
$\mathbb{E}\left[V_{1}(\chi_{\ell+1})+\bar{g}\left(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell}\right)\,\middle|\,\hat{\chi}_{\ell}\right]-\mathbb{E}\left[V_{1}(\chi_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right$] (119)
$=\hat{\chi}_{\ell}\left(P^{(\varphi(\hat{\chi}_{\ell}))}_{0}-P^{(1)}_{0}\right)\hat{\chi}_{\ell}+\operatorname{tr}\left(\left(P^{(\varphi(\hat{\chi}_{\ell}))}_{0}-P^{(1)}_{0}\right)\Sigma\right)$ (120)
$\quad+\beta^{(\varphi(\hat{\chi}_{\ell}))}+\gamma^{(\varphi(\hat{\chi}_{\ell}))}$ (121)
[…]
where $\beta^{(1)}$ and $\gamma^{(1)}$ are given by (109) and (110), respectively, and $f_{m}$ is defined in (66). Summing (122) over $\ell=0,1,\dots,L-1$ for $L\in\mathbb{N}$ yields
$\sum_{\ell=0}^{L-1}\mathbb{E}\left[V_{1}(\chi_{\ell+1})+\bar{g}\left(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell}\right)\,\middle|\,\hat{\chi}_{\ell}\right$] (123)
$\quad-\sum_{\ell=0}^{L-1}\mathbb{E}\left[V_{1}(\chi_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right$] (124)
[…]
Conditioning on $\hat{\chi}_{0}$ and applying both the tower property and the linearity of conditional expectation [22, Th. 5.1.2], the left-hand side of (125) becomes
$\mathbb{E}\Biggl[\sum_{\ell=0}^{L-1}\mathbb{E}\left[V_{1}(\chi_{\ell+1})+\bar{g}\left(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell}\right)\,\middle|\,\hat{\chi}_{\ell}\right$] (126)
$\quad-\sum_{\ell=0}^{L-1}\mathbb{E}\left[V_{1}(\chi_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right]\,\Big|\,\hat{\chi}_{0}\Biggr]$ (127)
$=\mathbb{E}\left[\sum_{\ell=0}^{L-1}\bar{g}\left(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell}\right)\,\middle|\,\hat{\chi}_{0}\right]$ (128)
$\quad+\mathbb{E}\left[V_{1}(\chi_{L})\,\middle|\,\hat{\chi}_{0}\right]-\mathbb{E}\left[V_{1}(\chi_{0})\,\middle|\,\hat{\chi}_{0}\right].$ (129)
Hence, we have
$\mathbb{E}\left[\sum_{\ell=0}^{L-1}\bar{g}\left(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell}\right)\,\middle|\,\hat{\chi}_{0}\right]$ (130)
$=L\left(\beta^{(1)}+\gamma^{(1)}\right)-\mathbb{E}\left[\sum_{\ell=0}^{L-1}f_{m}(\hat{\chi}_{\ell})\,\middle|\,\hat{\chi}_{0}\right]$ (131)
$\quad-\mathbb{E}\left[V_{1}(\chi_{L})\,\middle|\,\hat{\chi}_{0}\right]+\mathbb{E}\left[V_{1}(\chi_{0})\,\middle|\,\hat{\chi}_{0}\right].$ (132)
Dividing both sides by $hL$ and taking the limit superior as $L\to\infty$, the left-hand side of (132) is expressed as
$\limsup_{L\to\infty}\frac{1}{hL}\mathbb{E}\left[\sum_{\ell=0}^{L-1}\bar{g}\left(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell}\right)\,\middle|\,\hat{\chi}_{0}\right]$ (133)
$=J^{a}(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}}).$ (134)
On the right-hand side of (132), we obtain
$\limsup_{L\to\infty}\frac{1}{h}\left(\beta^{(1)}+\gamma^{(1)}\right)$ (135)
$=\frac{1}{p}\left(\operatorname{tr}\left(\tilde{P}_{p}D_{p}\Omega_{w}^{(p)}D_{p}^{\top}\right)+\operatorname{tr}\left(M_{p}\Sigma\right)+d^{a}(p)+\theta\right)$ (136)
$=J^{a}(\mu^{u,\text{per}},\mu^{\delta,\text{per}}),$ (137)
and
$\limsup_{L\to\infty}\frac{1}{hL}\mathbb{E}\left[V_{1}(\chi_{0})\,\middle|\,\hat{\chi}_{0}\right]=0.$ (138)
Combining (134), (137), and (138) gives
$J^{a}(\mu^{u,\text{ro}},\mu^{\delta,\text{ro}})=J^{a}(\mu^{u,\text{per}},\mu^{\delta,\text{per}})$ (139)
$\qquad-\limsup_{L\to\infty}\frac{1}{hL}\mathbb{E}\left[V_{1}(\chi_{L})\,\middle|\,\hat{\chi}_{0}\right]$ (140)
$\qquad-\limsup_{L\to\infty}\frac{1}{hL}\mathbb{E}\left[\sum_{\ell=0}^{L-1}f_{m}(\hat{\chi}_{\ell})\,\middle|\,\hat{\chi}_{0}\right].$ (141)
In the reminder, we show that the limit superiors in (141) converge to nonnegative points. From Lemma 1, it follows that
$\limsup_{L\to\infty}\frac{1}{hL}\mathbb{E}\left[V_{1}(\chi_{L})\,\middle|\,\hat{\chi}_{0}\right]=0.$ (142)
Moreover, for every initial condition $\hat{\chi}_{0}$, Theorem 2 ensures
$\lim_{L\to\infty}\mathbb{E}\left[\sum_{\ell=0}^{L-1}\tilde{f}_{m}(\hat{\chi}_{\ell})\,\middle|\,\hat{\chi}_{0}\right]=\int_{\mathbb{R}^{n_{x}}}\tilde{f}_{m}(\chi)\,\pi_{\hat{\chi}}(d\chi),$ (143)
where $\pi_{\hat{\chi}}$ is a unique invariant measure of the chain $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ [24, Th. 14.0.1]. Note that the existence of $\pi_{\hat{\chi}}$ is guaranteed by Theorem 2. The above integrand is strictly positive due to the open set irreducibility of $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ (see Appendix -F for details). Applying the Stolz–Cesàro theorem [29] to (143) yields
$\limsup_{L\to\infty}\frac{1}{hL}\mathbb{E}\left[\sum_{\ell=0}^{L-1}f_{m}(\hat{\chi}_{\ell})\,\middle|\,\hat{\chi}_{0}\right]=\tilde{c}_{m}+\frac{1}{h},$ (144)
where
$\tilde{c}_{m}\coloneqq\frac{1}{h}\int_{\mathbb{R}^{n_{x}}}\tilde{f}_{m}(\chi)\,\pi_{\hat{\chi}}(d\chi).$ (145)
Combining (141), (142), and (144) establishes (63). ∎
-D Proof of Lemma 1
First, we show that there exists a sufficiently small constant $a_{1}>0$ such that
$\mathbb{E}\left[\bar{g}(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right]\geq a_{1}\hat{\chi}_{\ell}^{\top}\hat{\chi}_{\ell},\quad\forall\ell\in\mathbb{Z}_{\geq 0}.$ (146)
Since $R\succ 0$ and $\theta>0$, it follows that
$\mathbb{E}\left[\bar{g}(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right]\geq\mathbb{E}\left[\sum_{\tau=0}^{h-1}x_{\ell h+\tau}^{\top}Qx_{\ell h+\tau}\,\middle|\,\hat{\chi}_{\ell}\right$]. (147)
From Assumption 2, we can set $a_{1}=\lambda_{\min}(Q)>0$, which satisfies (146).
Furthermore, for any constant $b_{1}$ such that $b_{1}>a_{1}>0$ and $b_{1}I_{n_{x}}-\tilde{P}_{p}\succ 0$, we obtain
$\mathbb{E}\left[V_{1}(\chi_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right]\leq b_{1}\hat{\chi}_{\ell}^{\top}\hat{\chi}_{\ell}+b_{1}\operatorname{tr}\left(\Sigma\right)$ (148)
from (64).
Hence, we have
$\mathbb{E}\left[V_{1}(\chi_{\ell+1})\,\middle|\,\hat{\chi}_{\ell}\right$] $\leq\beta^{(1)}+\gamma^{(1)}+\mathbb{E}\left[V_{1}(\chi_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right$] (149)
$\quad-a_{1}\hat{\chi}_{\ell}^{\top}\hat{\chi}_{\ell}$ (150)
$\leq d_{1}\mathbb{E}\left[V_{1}(\chi_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right]+c_{1},$ (151)
where
$d_{1}$ $=1-\frac{a_{1}}{b_{1}}<1,$ (152)
$c_{1}$ $=\beta^{(1)}+\gamma^{(1)}+a_{1}\operatorname{tr}\left(\Sigma\right).$ (153)
By iterating (151) for $\ell=0,1,\dots,L-1$, we conclude
$\mathbb{E}\left[V_{1}(\chi_{L})\,\middle|\,\hat{\chi}_{0}\right]\leq d_{1}^{L}\mathbb{E}\left[V_{1}(\chi_{0})\,\middle|\,\hat{\chi}_{0}\right]+\sum_{i=0}^{L-1}d_{1}^{i}c_{1},$ (154)
which shows that $\mathbb{E}\left[V_{1}(\chi_{L})\,\middle|\,\hat{\chi}_{0}\right]$ remains bounded as $L\to\infty$ since $d_{1}<1$. ∎
-E Proof of Lemma 2
[…]
$P(x,\mathcal{A})\coloneqq\mathbb{P}\left(\hat{\chi}_{\ell+1}\in\mathcal{A}\,\middle|\,\hat{\chi}_{\ell}=x\right).$ (155)
In the following, we show that $P$ is a transition probability kernel and that the process $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ has the Markov property [30, Def. 8.4.3], i.e.,
$\mathbb{P}\left(\hat{\chi}_{\ell+1}\in\mathcal{A}\,\middle|\,\mathscr{F}_{\ell}\right)=\mathbb{P}\left(\hat{\chi}_{\ell+1}\in\mathcal{A}\,\middle|\,\hat{\chi}_{\ell}\right)\quad\text{a.s.}$ (156)
For each $\varphi\colon\mathbb{R}^{n_{x}}\to\mathcal{M}$ and $m\in\mathcal{M}$, define
$\mathcal{C}_{m}\coloneqq\left\{x\in\mathbb{R}^{n_{x}}\colon\varphi(x)=m\right\}.$ (157)
We note that $\mathcal{C}_{m}$ is Borel measurable, which can be described as follows: Since $\varphi$ on $\mathcal{M}$ is determined by the minimization of finitely many continuous functions, each $\mathcal{C}_{m}$ can be expressed as a finite intersection of inverse images of Borel sets under continuous functions, i.e.,
$\mathcal{C}_{m}$ $=\left(\bigcap_{i<m}\left\{x\colon g_{m}(x)<g_{i}(x)\right\}\right)$ (158)
$\qquad\cap\left(\bigcap_{i>m}\left\{x\colon g_{m}(x)\leq g_{i}(x)\right\}\right)$ (159)
$=\left(\bigcap_{i<m}h_{i,m}^{-1}\left((0,\infty)\right)\right)\cap\left(\bigcap_{i>m}h_{i,m}^{-1}\left([0,\infty)\right)\right),$ (160)
where $g_{m}(x)\coloneqq x^{\top}P^{(m)}_{0}x+\operatorname{tr}(P^{(m)}_{0}\Sigma)+\beta^{(m)}+\gamma^{(m)}$ and $h_{i,m}(x)\coloneqq g_{i}(x)-g_{m}(x)$. This implies that $\mathcal{C}_{m}$ is Borel measurable since each inverse image is open and Borel measurable [27, Th. 4.8]. Therefore, for any $m\in\mathcal{M}$, we have $\varphi^{-1}(\{m\})=\mathcal{C}_{m}\in\mathscr{B}(\mathbb{R}^{n_{x}})$, which establishes the measurability of $\varphi$ [31, Th. 1.12].
Moreover, using the matrices $\Phi_{m}$ and $\Gamma_{m}$, we define
$\Psi(x,\omega)\coloneqq\sum_{m\in\mathcal{M}}\mathbb{I}_{\mathcal{C}_{m}}(x)\left(\Phi_{m}x+\Gamma_{m}\omega\right).$ (161)
Then, $\hat{\chi}_{\ell+1}=\Psi(\hat{\chi}_{\ell},\bar{\omega}_{\ell})$ for a given $\varphi$. For each $x$ and $\omega$, the mapping $(x,\omega)\mapsto\Phi_{m}x+\Gamma_{m}\omega$ is continuous, and hence measurable [31, Th. 1.7]. Since the indicator function is measurable, the mapping $(x,\omega)\mapsto\mathbb{I}_{\mathcal{C}_{m}}(x)(\Phi_{m}x+\Gamma_{m}\omega)$ is measurable. Therefore, $\Psi$ is also measurable for any $m\in\mathcal{M}$ [31, Th. 1.8].
We define the collection of noises $w_{k}$ and $v_{k}$ over $h$ steps as
$\bar{\mathcal{W}}_{\ell}\coloneqq\{w_{\ell h},\dots,w_{(\ell+1)h-1},v_{\ell h+1},\dots,v_{(\ell+1)h}\}$ (162)
for later use. By construction, $\bar{\omega}_{\ell}$ is a measurable function of $(e_{\ell h},\,\bar{\mathcal{W}}_{\ell})$, and hence $\sigma(\bar{\omega}_{\ell})\subseteq\sigma(e_{\ell h})\vee\sigma(\bar{\mathcal{W}}_{\ell})$ holds from [31, Th. 1.12]. Moreover, the independence of the noise processes yields $\sigma(\bar{\mathcal{W}}_{\ell})\perp\sigma(\mathcal{I}_{\ell h})$ [32, Lemma 4.6]. Since $e_{\ell h}=x_{\ell h}-\mathbb{E}[x_{\ell h}\mid\mathcal{I}_{\ell h}]$ under Gaussian noises implies that $e_{\ell h}$ is independent of $\mathcal{I}_{\ell h}$ [32, Lemma 14.1], we have
$\left(\sigma(e_{\ell h})\vee\sigma(\bar{\mathcal{W}}_{\ell})\right)\perp\sigma(\mathcal{I}_{\ell h}),$ (163)
which implies $\sigma(\bar{\omega}_{\ell})\perp\sigma(\mathcal{I}_{\ell h})$ [32, Cor. 4.7]. Since $\mathscr{F}_{\ell}\subseteq\sigma(\mathcal{I}_{\ell h})$, we obtain
$\sigma(\bar{\omega}_{\ell})\perp\mathscr{F}_{\ell}.$ (164)
[…]
$=\int_{\mathbb{R}^{n_{x}h}}\mathbb{I}_{\mathcal{A}}\left(\Psi(\hat{\chi}_{\ell},\omega)\right)\,\nu_{\bar{\omega}}(d\omega)\quad\text{a.s.}$ (166)
by [22, Ex. 5.1.5], where $\nu_{\bar{\omega}}$ denotes the distribution of $\bar{\omega}_{\ell}$. For a given $x\in\mathbb{R}^{n_{x}}$, combining (155), (164), and (166) yields
$P(x,\mathcal{A})=\int_{\mathbb{R}^{n_{x}h}}\mathbb{I}_{\mathcal{A}}\left(\Psi(x,\omega)\right)\,\nu_{\bar{\omega}}(d\omega)$ (167)
[…]
$\mathbb{E}\left[\mathbb{I}_{\mathcal{A}}\left(\Psi(\hat{\chi}_{\ell},\bar{\omega}_{\ell})\right)\,\middle|\,\sigma(\hat{\chi}_{\ell})\right]=\mathbb{E}\left[\mathbb{I}_{\mathcal{A}}\left(\Psi(\hat{\chi}_{\ell},\bar{\omega}_{\ell})\right)\,\middle|\,\mathscr{F}_{\ell}\right]\quad\text{a.s.},$ (168)
and hence, we obtain
$\mathbb{E}\left[\mathbb{I}_{\mathcal{A}}\left(\Psi(\hat{\chi}_{\ell},\bar{\omega}_{\ell})\right)\,\middle|\,\mathscr{F}_{\ell}\right]$ (169)
$=\int_{\mathbb{R}^{n_{x}h}}\mathbb{I}_{\mathcal{A}}\left(\Psi(\hat{\chi}_{\ell},\omega)\right)\,\nu_{\bar{\omega}}(d\omega)\quad\text{a.s.}$ (170)
from (166) and (168) [22, Th. 5.1.5].
For a fixed $x\in\mathbb{R}^{n_{x}}$, we define the mapping $\Psi_{x}\colon\omega\mapsto\Psi(x,\omega)$. We note that $\Psi_{x}$ is measurable for each $x$ since $\Psi$ is measurable on $\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{x}h}$ [31, Th. 8.5]. Then, for every $\mathcal{A}\in\mathscr{B}(\mathbb{R}^{n_{x}})$, it holds that
$P(x,\mathcal{A})$ $=\mathbb{P}\left(\hat{\chi}_{\ell+1}\in\mathcal{A}\,\middle|\,\hat{\chi}_{\ell}=x\right)=\mathbb{P}\left(\Psi(x,\bar{\omega}_{\ell})\in\mathcal{A}\right)$ (171)
$=\mathbb{P}\left(\Psi_{x}(\bar{\omega}_{\ell})\in\mathcal{A}\right)=\mathbb{E}\left[\mathbb{I}_{\mathcal{A}}\left(\Psi_{x}(\bar{\omega}_{\ell})\right)\right]$ (172)
$=\int_{\mathbb{R}^{n_{x}h}}\mathbb{I}_{\mathcal{A}}\left(\Psi_{x}(\omega)\right)\,\nu_{\bar{\omega}}(d\omega)$ (173)
$=\nu_{\bar{\omega}}\left(\left\{\omega\in\mathbb{R}^{n_{x}h}\colon\Psi_{x}(\omega)\in\mathcal{A}\right\}\right)$ (174)
$=\nu_{\bar{\omega}}\circ\Psi_{x}^{-1}(\mathcal{A}),$ (175)
which means that the mapping $\mathcal{A}\mapsto P(x,\mathcal{A})$ is a pushforward measure induced by $\Psi_{x}$; hence, $\mathcal{A}\mapsto P(x,\mathcal{A})$ is a probability measure on $\mathscr{B}(\mathbb{R}^{n_{x}})$ [32, Lemma 1.24]. Furthermore, $x\mapsto P(x,\mathcal{A})$ is measurable for a fixed $\mathcal{A}$, which follows by combining (167) and Fubini’s theorem. Therefore, $P(x,\mathcal{A})$ satisfies the definition of the transition probability kernel [24, Sec. 3.4].
By the property of conditional expectation [30, Th. 5.4.6], for any sub $\sigma$-algebra $\mathscr{G}\subseteq\mathscr{B}(\mathbb{R}^{n_{x}})$, we have
$\mathbb{P}\left(\hat{\chi}_{\ell+1}\in\mathcal{A}\,\middle|\,\mathscr{G}\right)=\mathbb{E}\left[\mathbb{I}_{\mathcal{A}}\left(\hat{\chi}_{\ell+1}\right)\,\middle|\,\mathscr{G}\right]\quad\text{a.s.}$ (176)
Hence, we obtain
$\mathbb{P}\left(\hat{\chi}_{\ell+1}\in\mathcal{A}\,\middle|\,\mathscr{F}_{\ell}\right)$ $=\int_{\mathbb{R}^{n_{x}h}}\mathbb{I}_{\mathcal{A}}\left(\Psi(\hat{\chi}_{\ell},\omega)\right)\,\nu_{\bar{\omega}}(d\omega)$ (177)
$=\mathbb{P}\left(\hat{\chi}_{\ell+1}\in\mathcal{A}\,\middle|\,\hat{\chi}_{\ell}\right)\quad\text{a.s.},$ (178)
which establishes (156). Therefore, the process $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ is a time-homogeneous Markov chain with the transition probability kernel given by (155). ∎
-F Proof of Theorem 2
In order to prove Theorem 2, we establish the following three lemmas:
Lemma 7
For any $x\in\mathbb{R}^{n_{x}}$ and any open set $\mathcal{O}\in\mathscr{B}(\mathbb{R}^{n_{x}})$, it holds that
$P^{n}(x,\mathcal{O})>0,\quad\forall n\geq 1,$ (179)
where $P$ is defined in (155).
Lemma 8
Let $\mathcal{B}_{\varepsilon}(z)\coloneqq\{x\in\mathbb{R}^{n_{x}}\colon\|z-x\|<\varepsilon\}$ denote the $\varepsilon$-neighborhood of $z$ for a constant $\varepsilon>0$. Then, for any $z\in\mathbb{R}^{n_{x}}$ and $\mathcal{A}\in\mathscr{B}(\mathbb{R}^{n_{x}})$, there exists a nonnegative continuous function $z\mapsto T(z,\mathcal{A})$ such that, for any $x\in\mathcal{B}_{\varepsilon}(z)$,
$P(x,\mathcal{A})\geq T(x,\mathcal{A}),\quad T(x,\mathbb{R}^{n_{x}})=1.$ (180)
Lemma 9
The function $V_{1}$ satisfies the following drift condition:
$\mathbb{E}\left[V_{1}(\hat{\chi}_{\ell+1})\,\middle|\,\hat{\chi}_{\ell}\right]-V_{1}(\hat{\chi}_{\ell})\leq-1,\quad\forall\hat{\chi}_{\ell}\in\mathbb{R}^{n_{x}}\backslash\mathcal{D},$ (181)
where
$\mathcal{D}\coloneqq\left\{\chi\in\mathbb{R}^{n_{x}}\colon a_{1}\chi^{\top}\chi\leq\beta^{(1)}+\gamma^{(1)}+1\right\}.$ (182)
First, we show Lemma 7. For each $m\in\mathcal{M}$, we define the function
$p_{m}(x,z)$ (183)
$\coloneqq\frac{\exp\left[-\tfrac{1}{2}(z-\Phi_{m}x)^{\top}\left(\Gamma_{m}\bar{\Omega}_{\omega}\Gamma_{m}^{\top}\right)^{-1}(z-\Phi_{m}x)\right]}{(2\pi)^{n_{x}/2}\sqrt{\det\!\left(\Gamma_{m}\bar{\Omega}_{\omega}\Gamma_{m}^{\top}\right)}},$ (184)
where $\bar{\Omega}_{\omega}$ is given by
$\bar{\Omega}_{\omega}$ $\coloneqq I_{h}\otimes\Omega_{\omega},$ (185)
$\Omega_{\omega}$ $\coloneqq G\big(CA\Sigma A^{\top}C^{\top}+C\Omega_{w}C^{\top}+\Omega_{v}\big)G^{\top}.$ (186)
Since $\Omega_{v}\succ 0$ and Assumption 2 holds, the matrix $\Omega_{\omega}$ is positive definite, and $\bar{\Omega}_{\omega}\succ 0$ also holds. Moreover, since $\Gamma_{m}$ has full row rank for every $m\in\mathcal{M}$ by its construction, it follows that $\Gamma_{m}\bar{\Omega}_{\omega}\Gamma_{m}^{\top}$ is positive definite. As $\bar{\omega}_{\ell}$ is Gaussian with covariance matrix $\bar{\Omega}_{\omega}$ and independent of $\hat{\chi}_{\ell}$, the state $\hat{\chi}_{\ell+1}$ conditioned on $\hat{\chi}_{\ell}=x$ and $\varphi(x)=m$ is Gaussian with mean $\Phi_{m}x$ and covariance matrix $\Gamma_{m}\bar{\Omega}_{\omega}\Gamma_{m}^{\top}\succ 0$. Therefore, the transition probability kernel $P$ admits a smooth density $P(x,dz)=p_{m}(x,z)\,dz$ for a fixed $x\in\mathbb{R}^{n_{x}}$ and $m=\varphi(x)$, where the mapping $z\mapsto p_{m}(x,z)$ is continuous and strictly positive on $\mathbb{R}^{n_{x}}$ [24, Sec. 4.4].
For any $x\in\mathbb{R}^{n_{x}}$ and $\mathcal{A}\in\mathscr{B}(\mathbb{R}^{n_{x}})$, we have
$P(x,\mathcal{A})$ $=\int_{\mathbb{R}^{n_{x}h}}\mathbb{I}_{\mathcal{A}}\left(\Psi(x,\omega)\right)\,\nu_{\bar{\omega}}(d\omega)$ (187)
$=\sum_{m\in\mathcal{M}}\mathbb{I}_{\mathcal{C}_{m}}(x)\int_{\mathbb{R}^{n_{x}h}}\mathbb{I}_{\mathcal{A}}\left(\Phi_{m}x+\Gamma_{m}\omega\right)\,\nu_{\bar{\omega}}(d\omega)$ (188)
$=\sum_{m\in\mathcal{M}}\mathbb{I}_{\mathcal{C}_{m}}(x)\int_{\mathcal{A}}p_{m}(x,z)\,dz,$ (189)
where, for each fixed $x\in\mathbb{R}^{n_{x}}$ and $m\in\mathcal{M}$, we use the relation
$\int_{\mathbb{R}^{n_{x}h}}\mathbb{I}_{\mathcal{A}}\left(\Phi_{m}x+\Gamma_{m}\omega\right)\,\nu_{\bar{\omega}}(d\omega)$ $=\nu_{\bar{\omega}}\circ\Psi_{x}^{-1}(\mathcal{A})$ (190)
$=\int_{\mathcal{A}}p_{m}(x,z)\,dz.$ (191)
The first equality is from (175). For the second equality, we note that $P(x,\cdot)=\nu_{\bar{\omega}}\circ\Psi_{x}^{-1}$ is absolutely continuous with respect to the Lebesgue measure and its Radon–Nikodym derivative is $p_{m}(x,\cdot)$. Hence, by the Radon–Nikodym theorem [31, Th. 6.10], it holds that
$\nu_{\bar{\omega}}\circ\Psi_{x}^{-1}(\mathcal{A})=\int_{\mathcal{A}}p_{m}(x,z)\,dz$ (192)
with $m=\varphi(x)$. Let $\mathcal{O}\subset\mathbb{R}^{n_{x}}$ be a nonempty open set. Since $p_{m}(x,\cdot)>0$ holds everywhere for any $m\in\mathcal{M}$, it follows that
$\int_{\mathcal{O}}p_{m}(x,z)\,dz>0.$ (193)
Thus, by (189), we have
$P(x,\mathcal{O})=\sum_{m\in\mathcal{M}}\mathbb{I}_{\mathcal{C}_{m}}(x)\int_{\mathcal{O}}p_{m}(x,z)\,dz>0$ (194)
for all $x\in\mathbb{R}^{n_{x}}$. By the Chapman–Kolmogorov equation with $P^{0}(x,\cdot)=\delta_{x}$ [24, Th. 3.4.2], for any $n\geq 1$, we obtain
$P^{n}(x,\mathcal{O})=\int_{\mathbb{R}^{n_{x}}}P^{n-1}(x,dy)P(y,\mathcal{O})>0,$ (195)
since $P(y,\mathcal{O})>0$ for all $y\in\mathbb{R}^{n_{x}}$ and $P^{n-1}(x,\cdot)$ is a probability measure on $\mathscr{B}(\mathbb{R}^{n_{x}})$[22, Ex. 1.4.1].
In the next step, we prove Lemma 8. For any $\mathcal{A}\in\mathscr{B}(\mathbb{R}^{n_{x}})$, we define
$T(x,\mathcal{A})\coloneqq\min_{m\in\mathcal{M}}\tilde{P}_{m}(x,\mathcal{A}),\quad\forall x\in\mathbb{R}^{n_{x}},$ (196)
where, for each $m\in\mathcal{M}$,
$\tilde{P}_{m}(x,\mathcal{A})$ $\coloneqq\mathbb{P}\left(\Phi_{m}x+\Gamma_{m}\bar{\omega}_{0}\in\mathcal{A}\right)$ (197)
$=\nu_{\bar{\omega}}\left(\{\omega\colon\Phi_{m}x+\Gamma_{m}\omega\in\mathcal{A}\}\right).$ (198)
Since $\mathcal{M}$ is finite, the minimum over $m\in\mathcal{M}$ always exists and (196) is well posed.
We first show that the mapping $x\mapsto T(x,\mathcal{A})$ is nonnegative. From (198), $\tilde{P}_{m}(x,\cdot)$ is the pushforward measure under the measurable map $\omega\mapsto\Phi_{m}x+\Gamma_{m}\omega$ for fixed $x\in\mathbb{R}^{n_{x}}$ and $m\in\mathcal{M}$. Hence, $\mathcal{A}\mapsto\tilde{P}_{m}(x,\mathcal{A})$ is a probability measure on $\mathscr{B}(\mathbb{R}^{n_{x}})$. Therefore, $\tilde{P}_{m}(x,\mathcal{A})\geq 0$ for all $m\in\mathcal{M}$, which gives $T(x,\mathcal{A})\geq 0$ for all $x\in\mathbb{R}^{n_{x}}$.
We then show that $x\mapsto T(x,\mathcal{A})$ is continuous. From (189), for any fixed $m\in\mathcal{M}$ we have
$\tilde{P}_{m}(x,\mathcal{A})=\int_{\mathcal{A}}p_{m}(x,z)\,dz.$ (199)
Let $\xi\in\mathbb{R}^{n_{x}}$ be fixed and set $\Delta_{\Phi}\coloneqq\Phi_{m}(x-\xi)$. We define the function
$r(y)\coloneqq\frac{\exp\left[-\tfrac{1}{2}\,y^{\top}\left(\Gamma_{m}\bar{\Omega}_{\omega}\Gamma_{m}^{\top}\right)^{-1}y\right]}{(2\pi)^{n_{x}/2}\sqrt{\det\left(\Gamma_{m}\bar{\Omega}_{\omega}\Gamma_{m}^{\top}\right)}},$ (200)
so that $p_{m}(x,z)=r(z-\Phi_{m}x)=r\bigl(z-\Phi_{m}\xi-\Delta_{\Phi}\bigr)$ and $p_{m}(\xi,z)=r(z-\Phi_{m}\xi)$. Then, we have
$\left|\tilde{P}_{m}(x,\mathcal{A})-\tilde{P}_{m}(\xi,\mathcal{A})\right|$ (201)
$\leq\int_{\mathcal{A}}\left|p_{m}(x,z)-p_{m}(\xi,z)\right|\,dz$ (202)
$\leq\int_{\mathbb{R}^{n_{x}}}\left|p_{m}(x,z)-p_{m}(\xi,z)\right|\,dz$ (203)
$=\int_{\mathbb{R}^{n_{x}}}\left|r(z-\Phi_{m}\xi-\Delta_{\Phi})-r(z-\Phi_{m}\xi)\right|\,dz$ (204)
$=\|\,r(\cdot-\Delta_{\Phi})-r(\cdot)\,\|_{L^{1}},$ (205)
where the first equality follows from the translation invariance of the Lebesgue measure [31, Th. 2.20]. Since $r\in L^{1}(\mathbb{R}^{n_{x}})$ and translations are continuous in $L^{1}$, it holds that
$\|\,r(\cdot-\Delta_{\Phi})-r(\cdot)\,\|_{L^{1}}\to 0\quad(\Delta_{\Phi}\to 0).$ (206)
by [31, Th. 9.5]. Hence, $\tilde{P}_{m}(x,\mathcal{A})$ is continuous in $x$ for each $m\in\mathcal{M}$. Since $\mathcal{M}$ is finite, the minimum of the finitely many continuous functions $x\mapsto\tilde{P}_{m}(x,A)$ is also continuous, which implies that $x\mapsto T(x,A)$ is continuous.
By the definition of $P$, for any $\varepsilon>0$ and all $x\in\mathcal{B}_{\varepsilon}(z)$, we have
$P(x,\mathcal{A})=\tilde{P}_{\varphi(x)}(x,\mathcal{A})\geq\min_{m\in\mathcal{M}}\tilde{P}_{m}(x,\mathcal{A})=T(x,\mathcal{A}).$ (207)
Moreover, since the map $\mathcal{A}\mapsto\tilde{P}_{m}(x,\mathcal{A})$ is a probability measure for every $m\in\mathcal{M}$, we obtain
$T(x,\mathbb{R}^{n_{x}})=\min_{m\in\mathcal{M}}\tilde{P}_{m}(x,\mathbb{R}^{n_{x}})=1,$ (208)
since $\tilde{P}_{m}(x,\mathbb{R}^{n_{x}})=1$ for all $m\in\mathcal{M}$ and $x\in\mathbb{R}^{n_{x}}$.
Afterwards, we prove Lemma 9. Since $\hat{x}_{k}$ is a sufficient statistics with respect to $\mathcal{I}_{k}$, and by the orthogonal property of the Kalman filter, it follows that, for any $k\in\mathbb{Z}_{\geq 0}$,
$\mathbb{E}\left[e_{k}^{\top}\hat{x}_{k}\,\middle|\,\hat{x}_{k}\right]$ $=\mathbb{E}\left[(x_{k}-\hat{x}_{k})^{\top}\hat{x}_{k}\,\middle|\,\hat{x}_{k}\right]$ (209)
$=\left(\mathbb{E}\left[x_{k}\,\middle|\,\hat{x}_{k}\right]-\hat{x}_{k}\right)^{\top}\hat{x}_{k}$ (210)
$=0.$ (211)
Hence, we obtain
$\mathbb{E}\left[V_{1}(\chi_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right$] $=\mathbb{E}\left[(\hat{\chi}_{\ell}+e_{\ell h})^{\top}P^{(1)}_{0}(\hat{\chi}_{\ell}+e_{\ell h})\,\middle|\,\hat{\chi}_{\ell}\right$] (212)
$=\mathbb{E}\left[V_{1}(\hat{\chi}_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right]+\operatorname{tr}\left(P^{(1)}_{0}\Sigma\right),$ (213)
and
$\mathbb{E}\left[V_{1}(\chi_{\ell+1})\,\middle|\,\hat{\chi}_{\ell}\right$] (214)
$=\mathbb{E}\left[(\hat{\chi}_{\ell+1}+e_{(\ell+1)h})^{\top}P^{(1)}_{0}(\hat{\chi}_{\ell+1}+e_{(\ell+1)h})\,\middle|\,\hat{\chi}_{\ell}\right$] (215)
$=\mathbb{E}\left[V_{1}(\hat{\chi}_{\ell+1})\,\middle|\,\hat{\chi}_{\ell}\right]+\operatorname{tr}\left(P^{(1)}_{0}\Sigma\right).$ (216)
Therefore, we have
$\mathbb{E}\left[V_{1}(\chi_{\ell+1})+\bar{g}\left(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell}\right)\,\middle|\,\hat{\chi}_{\ell}\right]-\mathbb{E}\left[V_{1}(\chi_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right$] (217)
$=\mathbb{E}\left[V_{1}(\hat{\chi}_{\ell+1})+\bar{g}\left(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell}\right)\,\middle|\,\hat{\chi}_{\ell}\right]-\mathbb{E}\left[V_{1}(\hat{\chi}_{\ell})\,\middle|\,\hat{\chi}_{\ell}\right$] (218)
[…]
where the second equality follows from (122). Using (146), we obtain
[…]
$=\beta^{(1)}+\gamma^{(1)}-f_{m}(\hat{\chi}_{\ell})-\mathbb{E}\left[\bar{g}\left(\chi_{\ell},\hat{\chi}_{\ell},\bar{w}_{\ell},\bar{\omega}_{\ell}\right)\,\middle|\,\hat{\chi}_{\ell}\right$] (221)
$\leq\beta^{(1)}+\gamma^{(1)}-a_{1}\hat{\chi}_{\ell}^{\top}\hat{\chi}_{\ell}$. (222)
From the definition of the set $\mathcal{D}$ in (182), it follows that, for any $\hat{\chi}_{\ell}\in\mathbb{R}^{n_{x}}\backslash\mathcal{D}$,
$\mathbb{E}\left[V_{1}(\hat{\chi}_{\ell+1})\,\middle|\,\hat{\chi}_{\ell}\right]-V_{1}(\hat{\chi}_{\ell})\leq-1.$ (223)
Building upon Lemmas 7, 8, and 9, we are ready to prove that the process $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ is $\tilde{f}_{m}$-ergodic. From Lemma 7, the process $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}{\geq 0}}$ is aperiodic and open set irreducible [24, Sec. 5.4, 6.1]. Furthermore, Lemma 8 shows that the chain is a T-chain. Hence, every compact set in $\mathbb{R}^{n_{x}}$ is petite [24, Th. 6.2.5], which yields that the set $\mathcal{D}$ is petite. Then, Lemmas 6 and 9 imply that the chain is non-evanescent [24, Th. 9.4.1], which is equivalent to Harris recurrence of $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}{\geq 0}}$ [24, Th. 9.2.2]. In addition, a Harris recurrent chain admits a unique invariant measure on $\mathscr{B}(\mathbb{R}^{n_{x}})$ [24, Th. 10.0.1]. Combining this fact with Lemma 9, we conclude that $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}{\geq 0}}$ is positive Harris recurrent with invariant measure $\pi_{\hat{\chi}}$ [24, Th. 11.0.1]. Moreover, it follows that $\tilde{f}_{m}(x)\geq 1$ for all $x\in\mathbb{R}^{n_{x}}$ by construction, which establishes the $\tilde{f}_{m}$-ergodicity of the chain $\{\hat{\chi}_{\ell}\}_{\ell\in\mathbb{Z}_{\geq 0}}$ [24, Th. 14.0.1]. ∎
-G Proof of Theorem 3
Since $P^{(1)}_{0}$ is positive definite by Lemma 4, from [23, Cor. 10.4.2], there exist constants $0<\underline{\lambda}\leq\overline{\lambda}<\infty$ such that
$\underline{\lambda}\|x\|^{2}\leq V_{1}(x)\leq\overline{\lambda}\|x\|^{2},\quad\forall x\in\mathbb{R}^{n_{x}}.$ (224)
By Lemma 1, there exist a constant $C_{0}<\infty$ and $\ell_{0}\in\mathbb{Z}_{\geq 0}$ such that
$\sup_{\ell\geq\ell_{0}}\,\mathbb{E}\left[V_{1}(\chi_{\ell})\right]\leq C_{0}.$ (225)
For the case where $0\leq\ell<\ell_{0}$, since the plant is linear and the noise process has bounded covariance, we have $\mathbb{E}\left[\|\chi_{\ell}\|^{2}\right]<\infty$, and hence, by (224), also $\mathbb{E}\left[V_{1}(\chi_{\ell})\right]<\infty$. Therefore,
$\sup_{\ell\geq 0}\mathbb{E}\left[V_{1}(\chi_{\ell})\right]<\infty.$ (226)
Combining (224) and (226) yields
$\sup_{\ell\geq 0}\mathbb{E}\left[\|\chi_{\ell}\|^{2}\right]\leq\frac{1}{\underline{\lambda}}\,\sup_{\ell\geq 0}\mathbb{E}\left[V_{1}(\chi_{\ell})\right]<\infty.$ (227)
For a fixed $m\in\mathcal{M}$, we have
$x_{k+1}$ $=Ax_{k}+Bu_{k}+w_{k}$ (228)
$=\Theta^{(m)}_{k-\ell h}x_{k}+\rho^{(m)}_{k-\ell h}BF^{(m)}_{k-\ell h}e_{k}+w_{k},$ (229)
for $k=\ell h,\ldots,(\ell+1)h-1$. Defining
$\Xi^{(m)}_{(\tau:0)}\coloneqq\Theta^{(m)}_{\tau-1}\cdots\Theta^{(m)}_{0},\quad\Xi^{(m)}_{(0:0)}\coloneqq I_{n_{x}},$ (230)
and unrolling (229) from $k=\ell h$ to $k=\ell h+\tau$ with $\tau\in\{0,1,\dots,h-1\}$, we obtain
$x_{\ell h+\tau}$ $=\Xi^{(m)}_{(\tau:0)}\chi_{\ell}$ (231)
$\quad+\sum_{i=0}^{\tau-1}\Xi^{(m)}_{(\tau:i+1)}\left(\rho^{(m)}_{i}BF^{(m)}_{i}e_{\ell h+i}+w_{\ell h+i}\right).$ (232)
Since $\Theta^{(m)}_{i}$ and $F^{(m)}_{i}$ are bounded for every $i\in\{0,1,\dots,h-1\}$ and $m\in\mathcal{M}$, we can define finite constants as
$\overline{\Theta}$ $\coloneqq\max_{m,i}\|\Theta^{(m)}_{i}\|,\quad\overline{F}\coloneqq\max_{m,i}\|BF^{(m)}_{i}\|$, (233)
$c_{\Theta}(\tau)$ $\coloneqq\max_{m}\|\Xi^{(m)}_{(\tau:0)}\|\leq\overline{\Theta}^{\tau}.$ (234)
From the Cauchy–Schwarz inequality, which gives $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$ for scalars $a$, $b$, and $c$, we obtain
$\mathbb{E}\left[\|x_{\ell h+\tau}\|^{2}\right]$ (235)
$\leq 3\mathbb{E}\left[\|\Xi^{(m)}_{(\tau:0)}\chi_{\ell}\|^{2}\right]$ (236)
$\quad+3\sum_{i=0}^{\tau-1}\mathbb{E}\left[\|\Xi^{(m)}_{(\tau:i+1)}\rho^{(m)}_{i}BF^{(m)}_{i}e_{\ell h+i}\|^{2}\right]$ (237)
$\quad+3\sum_{i=0}^{\tau-1}\mathbb{E}\left[\|\Xi^{(m)}_{(\tau:i+1)}w_{\ell h+i}\|^{2}\right]$ (238)
$\leq 3c_{\Theta}(\tau)^{2}\,\mathbb{E}\left[\|\chi_{\ell}\|^{2}\right]$ (239)
$\quad+3\sum_{i=0}^{\tau-1}\overline{\Theta}^{2(\tau-i-1)}\overline{F}^{2}\operatorname{tr}(\Sigma)$ (240)
$\quad+3\sum_{i=0}^{\tau-1}\overline{\Theta}^{2(\tau-i-1)}\operatorname{tr}\left(\Omega_{w}\right)$ (241)
$\leq a_{\tau}\mathbb{E}\left[\|\chi_{\ell}\|^{2}\right]+b_{\tau},$ (242)
where $a_{\tau}$ and $b_{\tau}$ are defined as
$a_{\tau}$ $\coloneqq 3\overline{\Theta}^{2\tau},$ (243)
$b_{\tau}$ $\coloneqq 3\sum_{i=0}^{\tau-1}\overline{\Theta}^{2(\tau-i-1)}\left(\overline{F}^{2}\operatorname{tr}(\Sigma)+\operatorname{tr}(\Omega_{w})\right).$ (244)
Taking the supremum in $\ell$ in (242) and using (227) yields
$\sup_{\ell\geq 0}\mathbb{E}\left[\|x_{\ell h+\tau}\|^{2}\right]\leq a_{\tau}\sup_{\ell\geq 0}\mathbb{E}\left[\|\chi_{\ell}\|^{2}\right]+b_{\tau}<\infty$ (245)
for $\tau=0,1,\dots,h-1$, which establishes (67). ∎